### v1

# Total Gluon Helicity Contribution to Proton Spin from Lattice QCD

Abstract
We report a state-of-the-art lattice QCD calculation of the total gluon helicity contribution to proton spin, $\Delta G$. The calculation is done on ensembles at three different lattice spacings $a=\{0.08,0.09,0.11\}$ fm. By employing distillation + momentum smearing for proton external states, we extract the bare matrix elements of the topological current $K^{\mu}$ under the 5-HYP smeared Coulomb gauge fixing configurations. Furthermore, we apply a non-perturbative $\mathrm{RI/MOM}$ renormalization scheme augmented with the Cluster Decomposition Error Reduction (CDER) technique to determine the renormalization constants of $K^{\mu}$. The results obtained from different components $K^{t,i}$ (with $i$ being the direction of proton momentum or polarization) are consistent with Lorentz covariance within uncertainties. After extrapolating to the continuum limit, $\Delta G$ is found to be $\Delta G=0.231(17)^{\mathrm{sta.}}(33)^{\mathrm{sym.}}$ at the $\overline{\mathrm{MS}}$ scale ${\mu}^{2}=10\ \mathrm{GeV}^{2}$, which constitutes approximately $46(7)\%$ of the proton spin.
Introduction: Since the European Muon Collaboration (EMC) measured the quark spin contribution to proton spin through deep inelastic scattering with polarized muon and proton 1988364, the origin of proton spin has remained one of the most profound puzzles in modern particle and nuclear physics Ji:2020ena. Our current understanding is that the quark spin contribution accounts for only $\sim 30\%$ of the proton spin deFlorian:2009vb; Nocera:2014gqa; COMPASS:2015mhb; Alexandrou:2017oeh; Lin:2018obj; Liang:2018pis. The remaining contributions come from quark orbital angular momentum (OAM), as well as gluon spin and OAM. Two popular spin sum rules have been proposed in the literature to understand the proton spin structure. One is the Jaffe-Manohar sum rule Jaffe:1989jz which offers a complete decomposition of the proton spin into quark and gluon spin and OAM. It has a clear partonic picture in the infinite momentum frame (IMF) of the proton and lightcone gauge. The other is the frame- and gauge-independent sum rule by Ji Ji:1996ek, where the quark and gluon total angular momentum can be connected to the moments of generalized parton distributions Muller:1994ses; Ji:1996ek.
Among these contributions, the understanding of gluon spin has posed a significant challenge, mainly due to the absence of a local gauge-invariant gluon spin operator Ji:2012gc. Nevertheless, it can be measured in polarized scattering experiments through the spin-dependent gluon helicity distribution $\Delta g(x)$ as its lowest moment $\Delta G=\int dx\Delta g(x)$. Note that unlike in the case of the quark distribution, the lowest moment of the gluon helicity distribution is still nonlocal. Only in the lightcone gauge does it reduce to the matrix element of a local operator $\vec{E}\times\vec{A}$. Phenomenological global analyses of experimental data indicate that $\Delta G$ is approximately 0.2 at the $\overline{\mathrm{MS}}$ scale $\mu^{2}=10\ \mathrm{GeV}^{2}$, provided that contributions from the small-$x$ region ($x<0.05$) with large uncertainties are excluded deFlorian:2009vb; deFlorian:2014yva. On the other hand, recent theoretical developments Ji:2013fga; Hatta:2013gta; Ji:2014lra, particularly within the framework of Large-Momentum Effective Theory (LaMET) Ji:2013dva; Ji:2014gla; Ji:2020ect, have enabled a direct computation of $\Delta G$ using lattice QCD. In this approach, instead of computing in the IMF and lightcone gauge which are both inaccessible on the lattice, one computes the large-momentum nucleon matrix element of a static gluon spin operator, $\vec{E}\times\vec{A}$, in a physical gauge that leaves the transverse gauge field intact. Actually, there exists an infinite number of operators that can be used to extract $\Delta G$, all belonging to the same universality class Hatta:2013gta. The static gluon spin matrix element can then be matched to $\Delta G$ through a perturbative factorization or matching relation Ji:2013fga. Based on this, the $\rm\chi QCD$ collaboration conducted an exploratory calculation of $\Delta G$ by evaluating the nucleon matrix element of $\vec{E}\times\vec{A}$ in the Coulomb gauge, yielding a result of $0.251(47)(16)$ Yang:2016plb. However, the one-loop lattice perturbative renormalization employed in that study introduces a large correction, indicating the necessity of a proper nonperturbative renormalization – which was, at the time, very challenging to implement due to the Coulomb gauge fixing. Furthermore, recent studies Pang:2024sdl have revealed that the LaMET factorization for $\Delta G$ is incompatible with that for $\Delta g(x)$.
To circumvent the difficulties mentioned above, two new proposals have been made in Ref. Pang:2024sdl. One is based on using the topological current, and the other employs appropriate static nonlocal gluon operators generating the gluon helicity distribution. Both of them have a trivial matching only, and therefore resolve the inconsistency previously encountered in the matching for $\Delta G$ and $\Delta g(x)$. Moreover, their nonperturbative renormalization can be easily implemented. This opens a new avenue for systematic studies of $\Delta G$ from lattice QCD.
In this work, we present a state-of-the-art lattice calculation of the total gluon helicity contribution to proton spin, $\Delta G$, using the topological current approach proposed in Ref. Pang:2024sdl. The calculation is done on ensembles at three different lattice spacings $a=\{0.08,0.09,0.11\}$ fm, which allow us to perform a continuum extrapolation. Distillation HadronSpectrum:2009krc, momentum smearing Bali:2016lva, and cluster decomposition Liu:2017man techniques have been used to improve the signal-to-noise ratio of proton matrix elements and parton Green’s functions, respectively. The nonperturbative renormalization has been performed in the $\mathrm{RI/MOM}$ scheme, following the operator mixing pattern given in Ref. Pang:2024sdl. After converting to the $\overline{\rm MS}$ scheme, the result is then evolved to the scale $\mu^{2}=10\,{\rm GeV}^{2}$ using the renormalization group (RG) equation with three-loop anomalous dimensions.
Theoretical Framework: Following Ref. Pang:2024sdl, the total gluon helicity $\Delta G$ can be extracted by computing the matrix element of the local topological current
$K^{\mu}(x)=\epsilon^{\mu\nu\rho\sigma}\text{Tr}[A_{\nu}F_{\rho\sigma}-\frac{2}{3}ig_{0}A_{\nu}A_{\rho}A_{\sigma}](x)$ (1)
in a longitudinally polarized proton. More specifically, the proton matrix element of the Coulomb gauge fixed operator $K^{\mu}$ ($\mu=t,i$ with $i$ denoting the component along the direction of proton momentum or polarization) reduces to $\Delta G$ in the IMF, namely,
$\frac{\langle\mathrm{PS}_{\mathrm{Proj.i}}|\int d^{3}xK^{t/i}_{\mathrm{C.G.}}(x)|\mathrm{PS}_{\mathrm{Proj.i}}\rangle}{2S^{t/i}}|_{\mathrm{IMF}}=\Delta G,$ (2)
where $|\mathrm{PS}_{\mathrm{Proj.i}}\rangle$ denotes the proton external state with specific momentum and polarization direction $i$, while $S^{t/i}$ represents the corresponding polarization vector. The subscript $\mathrm{C.G.}$ indicates the Coulomb gauge fixing for the topological current.
As shown in Ref. Pang:2024sdl, the ultraviolet (UV) divergence of $K^{\mu}$ is gauge independent. Therefore, we can perform a nonperturbative renormalization of $K^{\mu}$ in the $\mathrm{RI/MOM}$ scheme MARTINELLI199581 where the renormalization factor is typically computed in the Landau gauge. $K^{\mu}$ mixes only with the axial-vector current $J^{\mu}_{5}=\bar{q}_{i}\gamma_{5}\gamma^{\mu}q_{i}$ even under RI/MOM renormalization Pang:2024sdl. The renormalization constants can be determined by the following conditions
$\begin{pmatrix}\langle ps|K^{\mu}|ps\rangle^{\mathrm{tree}}\\ \langle ps|J^{\mu}_{5}|ps\rangle^{\mathrm{tree}}\end{pmatrix}=\begin{pmatrix}Z_{11}^{\mathrm{RI}}&Z_{12}^{\mathrm{RI}}\\ Z_{21}^{\mathrm{RI}}&Z_{22}^{\mathrm{RI}}\end{pmatrix}\begin{pmatrix}\langle ps|K^{\mu}|ps\rangle^{\mathrm{lat.}}\\ \langle ps|J^{\mu}_{5}|ps\rangle^{\mathrm{lat.}}\end{pmatrix}$, (3)
where $|ps\rangle$ denotes either gluon ($g$) or quark ($q$) external states with specific momentum and polarization. In the discussion below, we introduce the abbreviations $\mathcal{O}^{\mathrm{tree/lat.},{g/q}}\equiv\langle{g,q}|\mathcal{O}|{g,q}\rangle^{\mathrm{tree/lat.}}$ to simplify the notation.
Note that $K^{\mathrm{tree},q}=J_{5}^{\mathrm{tree},g}=0$. Moreover, lattice data indicate that $J_{5}^{\mathrm{lat.},g}\approx 0$ within statistical uncertainties. Thus, the renormalization constants can be determined as:
$Z_{11}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\frac{K^{\mathrm{tree,g}}}{K^{\mathrm{lat.,g}}},\quad\quad Z_{12}^{\mathrm{RI}}(\mu_{\rm RI}^{2})=-\frac{K^{\mathrm{lat.,q}}}{J_{5}^{\mathrm{lat.,q}}}Z_{11}^{\mathrm{RI}},$
$Z_{21}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $0,\quad\quad\quad\quad\quad Z_{22}^{\mathrm{RI}}(\mu_{\rm RI}^{2})=\frac{J_{5}^{\mathrm{tree,q}}}{J_{5}^{\mathrm{lat.,q}}},$ (4)
where $\mu_{\rm RI}$ denotes the renormalization scale in the RI/MOM scheme. With the above approximation, $Z_{22}^{\mathrm{RI}}$ reduces to the axial-vector current renormalization constant $Z_{A}^{\mathrm{RI}}$ in Ref. Hu:2023jet. The ratio $K^{\mathrm{lat.,q}}/J_{5}^{\mathrm{lat.,q}}$ appearing in the off-diagonal element $Z_{12}^{\mathrm{RI}}$ is of $\mathcal{O}(\alpha_{s})$. Thus, the $\mathrm{RI/MOM}$-renormalized gluon and quark helicities are given by:
$\Delta G^{\mathrm{RI}}$ $=$ $Z_{11}^{\mathrm{RI}}\big(\Delta G^{\mathrm{B.}}-\frac{K^{\mathrm{lat.,q}}}{J_{5}^{\mathrm{lat.,q}}}\Delta\Sigma^{\mathrm{B.}}\big),$
$\Delta\Sigma^{\mathrm{RI}}$ $=$ $Z_{A}^{\mathrm{RI}}\Delta\Sigma^{\mathrm{B.}},$ (5)
where for simplicity the scale dependence has been suppressed. $\Delta G^{\mathrm{B.}}$ and $\Delta\Sigma^{\mathrm{B.}}$ denote the bare gluon and quark helicities extracted from the bare matrix elements of $K^{\mu}$ and $J_{5}^{\mu}$, respectively.
The RI/MOM renormalized results can then be converted to the $\overline{\mathrm{MS}}$ scheme via Pang:2024sdl
$\Delta G^{\overline{\mathrm{MS}}}$ $=$ $R_{11}\Delta G^{\mathrm{RI}}+R_{12}\Delta\Sigma^{\mathrm{RI}},$
$\Delta\Sigma^{\overline{\mathrm{MS}}}$ $=$ $R_{21}\Delta G^{\mathrm{RI}}+R_{22}\Delta\Sigma^{\mathrm{RI}},$ (6)
where again the scale dependence has been suppressed for simplicity. In the present work, we have included the conversion factors up to three-loops. Their detailed expressions are given in the Supplemental Material supplemental.
Lattice Calculation: Our calculation is carried out on the $n_{f}=$2+1 flavor full QCD ensembles generated by CLQCD using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action Hu:2023jet. The ensembles used for the simulation are described in Table 1. The values of $m_{\pi}$ of all the ensembles are chosen to ensure $m_{\pi}L\geq 3.7$, thereby efficiently suppressing the finite-volume effect. For more details, we refer the readers to the Supplemental Material supplemental.
Symbol $\hat{\beta}$ $a$ (fm) $\tilde{L}^{3}\times\tilde{T}$ $m_{\pi}$ (MeV) $N_{\mathrm{cfg}}$
C24P29 6.200 0.10524(05)(62) $24^{3}\times 72$ 292.3(1.0) 780
C48P23 6.200 0.10524(05)(62) $48^{3}\times 96$ 224.1(1.2) 400
E32P29 6.308 0.08973(20)(53) $32^{3}\times 64$ 287.3(2.5) 890
F32P30 6.410 0.07753(03)(45) $32^{3}\times 96$ 300.4(1.2) 800
Table 1: Gauge coupling $\hat{\beta}=10/(g_{0}^{2}u_{0}^{4})$, lattice spacing $a$, dimensionless lattice size $\tilde{L}^{3}\times\tilde{T}$, corresponding pion mass $m_{\pi}$, and number of configurations $N_{\mathrm{cfg}}$ used for the simulation.
In applying Eq. (2), we choose the proton momentum and polarization to be along three spatial directions and average over them, in order to enhance statistical precision while maintaining rotational symmetry. Furthermore, 5-HYP smearing is applied to the topological current throughout this calculation.
For proton external states at low momenta ($p\leq 0.5\ \mathrm{GeV}$), we implement the distillation smearing technique HadronSpectrum:2009krc, while for higher momentum states we employ an approach combining distillation with momentum smearing Bali:2016lva. We insert the topological current operator into the two-point function of a proton polarized along the $i$th-direction (denoted as $\mathrm{Proj.i}$) to obtain the disconnected three-point function
$C_{3}(t_{f},t_{i},t_{s})\equiv\langle\mathrm{PS}_{\mathrm{Proj.i}}(t_{f})|K^{t/i}_{\mathrm{C.G.}}(t_{i})|\mathrm{PS}_{\mathrm{Proj.i}}(t_{s})\rangle,$ (7)
where $t_{f}$, $t_{i}$, and $t_{s}$ correspond to sink, current insertion, and source time-slices, respectively. The bare matrix elements (BMEs) of the topological current are extracted through the following ratio:
$R_{K^{\mu}}(t_{f},t_{i},t_{s})=\frac{C_{3}(t_{f},t_{i},t_{s})}{C_{2}(t_{f},t_{s})},$ (8)
with $C_{2}(t_{f},t_{s})\equiv\langle\mathrm{PS}_{\mathrm{UnProj.}}(t_{f})|\mathrm{PS}_{\mathrm{UnProj.}}(t_{s})\rangle$ representing the unpolarized proton two-point correlation function. To optimize signal quality, we perform a summation over all source time-slices: $R_{K^{\mu}}(t_{f},t_{i})=\sum_{t_{s}}R_{K^{\mu}}(t_{f},t_{i},t_{s})$, which is then fitted to the form
$R_{K^{\mu}}(t_{f},t_{i})=\langle K^{\mu}\rangle+c_{1}e^{-\Delta E(t_{f}-t_{i})}+c_{2}e^{-\Delta Et_{i}},$ (9)
where $\langle K^{\mu}\rangle$ serves as an abbreviation for $\langle\mathrm{PS}_{\mathrm{Proj.i}}|K^{\mu}|\mathrm{PS}_{\mathrm{Proj.i}}\rangle/(2E)$, and $c_{1},c_{2},\Delta E$ are fitting parameters. For illustrative purposes, we show in FIG. 1 the fitted ground state matrix element for both $K^{i}$ and $K^{t}$ at $p=0.98$ GeV. The fitting results for other momenta can be found in the Supplemental Material supplemental.
Refer to caption Refer to caption
Figure 1: The ratio $R_{K^{\mu}}(t_{f},t_{i})$ at $t_{f}\in[0.63,1.16]\ \mathrm{fm}$ and the fitted BME $\langle K^{\mu}\rangle$ (the gray error band) for proton external states with p=0.98GeVp=0.98\ $\mathrm${GeV} under ensemble C24P29, utilizing the fitting procedure described in (9).
To obtain the renormalization constants, we evaluate the bare Green’s function for the topological current
$G_{K}^{\mu\rho\nu}(p_{1},p_{2})=\sum_{x,y}e^{-i(p_{1}\cdot x-p_{2}\cdot y)}\langle A^{\mu}(x)K^{\rho}(0)A^{\nu}(y)\rangle.$ (10)
The bare gluon propagator is given by
$S_{g}^{\mu\nu}(p)=\mathrm{CV}\langle\mathrm{Tr}[A^{\mu}(p)A^{\nu}(-p)]\rangle,$ (11)
where $\mathrm{CV}=2/[(N^{2}_{c}-1)\tilde{V}$], with $N_{c}=3$ representing the number of colors and $\tilde{V}=\tilde{L}^{3}\times\tilde{T}$ denoting the dimensionless lattice volume. Due to rotational symmetry breaking on the lattice, we can define the gluon field renormalization constants through two distinct prescriptions:
$Z_{g,\mathrm{diag}}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\frac{\sum_{p_{\rho}=0}1}{\hat{p}^{2}\sum_{p_{\rho}=0}S_{g}^{\rho\rho}(p^{2})},$
$Z_{g,\mathrm{off}}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\frac{\sum_{p_{\rho}\neq 0,p_{\tau}\neq 0,\rho\neq\tau}1}{\hat{p}^{4}\sum_{p_{\rho}\neq 0,p_{\tau}\neq 0,\rho\neq\tau}\frac{S_{g}^{\rho\tau}(p^{2})}{\hat{p}_{\rho}\hat{p}_{\tau}}},$ (12)
where $\hat{p}_{\mu}=\frac{2}{a}\sin{\frac{p_{\mu}a}{2}}$ represents the lattice momentum, and $\mu_{\rm RI}^{2}=p^{2}$ denotes the RI/MOM renormalization scale.
$Z_{11}^{\mathrm{RI}}$ in Eq.(Total Gluon Helicity Contribution to Proton Spin from Lattice QCD) then takes the following form
$Z_{11}^{\mathrm{RI}}(\mu_{\rm RI}^{2})=\frac{Z_{g,\mathrm{diag/off}}(\mu_{\rm RI}^{2})}{\mathrm{Tr}[S_{g}^{-1}(p_{1})G_{K}(p_{1},p_{2})S_{g}^{-1}(p_{2})(K^{\mathrm{tree,g}})^{-1}]},$ (13)
where the external momenta are chosen to satisfy $p_{1}=p_{2}=p$. The Green’s function of $K^{\mathrm{tree,g}}$ reads $i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}$ in our normalization convention. In addition, we hide the indicator of the bare Green’s function $G_{K}$ and bare gluon propagator $S_{g}$ in (13).
The inverse of $Z_{11}^{\rm RI}$ then reads,
$Z_{11}^{-1,\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\sum_{\mu,\rho,\nu,\sigma}\frac{2\mathrm{CV}\hat{p}^{2}G_{K}^{\mu\rho\nu}(p,p)}{i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}(S_{g}^{\mu\mu}+S_{g}^{\nu\nu})(p^{2})}|_{\mu_{\rm RI}^{2}=p^{2}}$ (14)
$=$ $\sum_{\mu,\rho,\nu,\sigma}\frac{\mathrm{CV}\hat{p}^{4}G_{K}^{\mu\rho\nu}(p,p)}{i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}\hat{p}_{\mu}\hat{p}_{\nu}S_{g}^{\mu\nu}(p^{2})}|_{\mu_{\rm RI}^{2}=p^{2}}.$
The summation over indices $\mu,\nu,\rho,\sigma$ in conjunction with the Levi-Civita antisymmetric tensor yields a total of 12 distinct configurations. To enhance the signal quality of the pure gluon Green’s function $G_{K}(p,p)$ in (14), we employ the cluster decomposition error reduction (CDER) technique Liu:2017man, which requires large volume ensembles. Due to the limited volume of C24P29, we have chosen to compute the renormalization constants on a larger volume ensemble, C48P23, which shares the same lattice spacing. In contrast, for the other two ensembles E32P29 and F32P30, the CDER can be directly applied. For more details on the CDER implementation, we refer the readers to the Supplemental Material supplemental.
The conversion factors from the RI/MOM scheme to the $\overline{\rm MS}$ scheme are evaluated by choosing $\mu/\mu_{\rm RI}=e^{-{5}/{6}}$ Brodsky:1982gc in the fixed-order expression up to three-loops, and then RG-evolved to $\mu^{2}=10\,{\rm GeV}^{2}$. Subsequently, the discretization effects of ${\cal O}(a^{2}p^{2})$ are removed following the strategy proposed in Ref. Yang:2018nqn. Note that here the extrapolation is conducted by varying $p^{2}$ (and thus $a^{2}p^{2}$) and determining the value at $a^{2}p^{2}=0$. To simultaneously ensure the reliability of RI/MOM renormalization and an acceptable signal-to-noise ratio after applying the CDER, we choose the data in the momentum interval $\mu_{\mathrm{RI}}\in[3.0,5.0]\ \mathrm{GeV}$ for the extrapolation, which correspond to $a^{2}p^{2}\in[2.6,7.1]$ under ensemble C48P23. The extrapolated result for the inverse of $\bar{Z}_{11}^{\mathrm{lat}}\equiv R_{11}Z_{11}^{\mathrm{RI}}$ is shown in Fig. 2. Similar analyses have also been performed for the other two ensembles with different lattice spacings.
Regarding the mixing with quark spin $\Delta\Sigma$, it turns out to be an order of magnitude less than $\bar{Z}_{11}^{\mathrm{lat}}$, especially at large $a^{2}p^{2}$ (See the Supplemental Material supplemental for more details). These properties offer a significant simplification in our numerical analysis, leading us to
$\Delta G^{\overline{\mathrm{MS}}}\approx\bar{Z}_{11}^{\mathrm{lat}}\Delta G^{\mathrm{B.}},\quad\Delta\Sigma^{\overline{\mathrm{MS}}}\approx\bar{Z}_{A}^{\mathrm{lat}}\Delta\Sigma^{\mathrm{B.}}.$ (15)
Figure 2: The removal of ${\cal O}(a^{2}p^{2})$ discretization effect in $\bar{Z}_{11}^{\mathrm{lat}}\equiv R_{11}Z_{11}^{\mathrm{RI}}$ for the ensemble C48P23. The red, green, and blue data points and their error bands represent the CDER results and their $a^{2}p^{2}\rightarrow 0$ extrapolated results using momentum modes $(0,0,p,p)$, $(0,p,p,p)$, and $(p,p,p,p)$ in (14), where $p=\big(2\pi/(aL)\big)p_{\mathrm{lat}}$ with $p_{\mathrm{lat}}$ represents momentum in the lattice unit.
Numerical Results: After obtaining the renormalized matrix elements of $K^{\mu}$ and converting them to the $\overline{\rm MS}$ scheme, we are able to extract the final result by carrying out a continuum and infinite momentum extrapolation. Fig. 3 displays $\langle K^{\mu}\rangle$ as a function of lattice spacings and proton momenta. As can be seen from the plot, the results for $\mu=t,i$ exhibit a clear trend consistent with Lorentz covariance within uncertainties. Motivated by this, we adopt the following form for the continuum and infinite momentum extrapolation:
$\langle K^{\mu}\rangle=\frac{S^{\mu}}{E}\big(\Delta G+\frac{c_{\mathrm{h.t.}}}{E^{2}}\big)+c_{a}a^{2},$ (16)
where $\Delta G$ is the result to be extracted, $c_{\mathrm{h.t.}}$ accounts for higher-twist contributions, and $c_{a}$ captures discretization effects. The pion mass dependence has been neglected in this analysis, primarily due to the significant computational cost of gluon matrix elements at the physical point and the very mild pion mass dependence observed in Ref. Yang:2016plb. Since the proton momenta are not significantly larger than its mass, we incorporate the higher-twist contribution using an empirical parametrization form $1/E^{2}$ in the equation above, where $E=\sqrt{M^{2}+p^{2}}$ with the proton mass $M\approx 1.02\ \mathrm{GeV}$ according to the dispersion relation. The primary sources of systematic uncertainty in our analysis are identified as follows: 1) Including or not the zero-momentum data points in the extrapolation; 2) Discretization errors associated with the finite lattice spacing; 3) The choice of the UV window in the RI/MOM scale or $a^{2}p^{2}$ (the interval $\mu_{\mathrm{RI}}\in[3.0,5.0]\ \mathrm{GeV}$ has been extended to [3.0,6.0]GeV[3.0,6.0]\ $\mathrm${GeV} to estimate the corresponding uncertainty); and 4) Systematic errors associated with the perturbative inputs (estimated by varying from three-loop to two-loop accuracy). As illustrated in FIG. 3, the joint extrapolation yields a gluon helicity contribution $\Delta G^{\overline{\mathrm{MS}},\ 10\ \mathrm{GeV}^{2}}=0.231(17)^{\mathrm{sta.}}(33)^{\mathrm{sym.}}$, which constitutes $46(7)\%$ of the proton’s total spin.
Figure 3: The continuum and infinite momentum extrapolation of the total gluon helicity results using lattice spacings ranging from 0.08fm0.08\ $\mathrm${fm} to 0.11fm0.11\ $\mathrm${fm} at the $\overline{\mathrm{MS}}$ scale $\mu^{2}=10\ \mathrm{GeV}^{2}$. The uncertainties are shown as separate bands: statistical uncertainty (dark gray) and all uncertainties that arise during our analysis (light gray).
Summary: In this work, we presented a state-of-the-art lattice QCD calculation of the total gluon helicity contribution, $\Delta G$, to the proton spin. The calculation leveraged appropriate components of the topological current operator, $K^{\mu}$ (specifically $\mu=\{t,i\}$ where $i$ is along the direction of the proton momentum). The bare matrix element is nonperturbatively renormalized in the RI/MOM scheme and subsequently converted to the $\overline{\rm MS}$ scheme via a three-loop conversion factor with a RG improvement. The result is then extrapolated to the continuum and infinite momentum limits, yielding $\Delta G=0.231(17)^{\mathrm{sta.}}(33)^{\mathrm{sym.}}$ at $\mu^{2}=10\ \mathrm{GeV}^{2}$. This corresponds to approximately $46(7)\%$ of the proton spin and clearly rules out the possibility of a negative gluon helicity contribution Zhou:2022wzm; Whitehill:2022mpq; Hunt-Smith:2024khs. The finding offers important insight for future theoretical and experimental efforts aimed at resolving the proton spin puzzle. We have also taken into account various sources of systematic uncertainties, including those from the exclusion of zero-momentum points, discretization errors, and the choice of the renormalization scale.
Moreover, our analysis demonstrated, for the first time, that the temporal and spatial components of $K^{\mu}$ generate consistent results for $\Delta G$ when the proton is boosted to the IMF. This is physically intuitive, as these two components are predominantly sensitive to color electric and magnetic fields, respectively, which become of comparable magnitude in the IMF.
Our calculation showcases the capability of combining high-precision extraction of disconnected matrix elements with advanced renormalization techniques to address challenging problems in lattice QCD. The result offers important guidance for future experimental measurements of gluon spin and paves the way for subsequent calculations of quark and gluon OAM, ultimately enabling a complete first-principles decomposition of the proton spin.
Acknowledgment
We thank Fangcheng He, Zhi-Cheng Hu, Xiangyu Jiang, Li-Jun Zhou, and other CLQCD and LPC members for valuable comments and suggestions. We thank the CLQCD collaborations for providing us their gauge configurations with dynamical fermions Hu:2023jet, which are generated on HPC Cluster of ITP-CAS, IHEP-CAS and CSNS-CAS, the Southern Nuclear Science Computing Center(SNSC) and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University. The calculations were performed using the PyQUDA Jiang:2024lto and Chroma Edwards:2004sx software suite with QUDA Clark:2009wm; Babich:2011np; Clark:2016rdz through HIP programming model Bi:2020wpt. The numerical calculation were carried out on the ORISE Supercomputer, the Southern Nuclear Science Computing Center (SNSC). This work is supported in part by NSFC grants No. 12375080, 12525504, 12435002, 12293060, 12293062, and 12447101, No. 12205171, No. 12235008, No. 12321005, National Key R&D Program of China No.2024YFE0109800, the Ministry of Science and Technology of China under Grant No. 2024YFA1611004, by CUHK-Shenzhen under grant No. UDF01002851, the Strategic Priority Research Program of Chinese Academy of Sciences, Grant No. YSBR-101, and Department of Science and Technology of Shandong province No. tsqn202312052 and 2024HWYQ-005.
Supplemental Material
.1 Detailed information of CLQCD ensembles
The results in this work, are based on the $n_{f}=2+1$ flavor ensembles from the CLQCD collaboration using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action Hu:2023jet.
The TITLS gauge action is defined in the following,
$S_{g}=\frac{1}{N_{c}}\mathrm{Re}\sum_{x,\mu<\nu}\mathrm{Tr}\left[1-\hat{\beta}\left(\mathcal{P}^{U}_{\mu,\nu}(x)+\frac{c_{1}\mathcal{R}^{U}_{\mu,\nu}(x)}{1-8c_{1}^{0}}\right)\right],$ (17)
and
$\mathcal{P}^{U}_{\mu,\nu}(x)$ $=U_{\mu}(x)U_{\nu}(x+a\hat{\mu})U^{\dagger}_{\mu}(x+a\hat{\nu})U^{\dagger}_{\nu}(x),$
$\mathcal{R}^{U}_{\mu,\nu}(x)$ $=U_{\mu}(x)U_{\mu}(x+a\hat{\mu})U_{\nu}(x+2a\hat{\mu})U^{\dagger}_{\mu}(x+a\hat{\mu}+a\hat{\nu})U^{\dagger}_{\mu}(x+a\hat{\nu})U^{\dagger}_{\nu}(x),$
$U_{\mu}(x)$ $=P\left[\mathrm{exp}\left({\rm i}g_{0}\int_{x}^{x+\hat{\mu}a}\mathrm{d}yA_{\mu}(y)\right)\right],$
$\hat{\beta}=(1-8c_{1}^{0})\frac{6}{g_{0}^{2}u_{0}^{4}}\equiv 10/(g_{0}^{2}u_{0}^{4})$ with $c_{1}^{0}=-\frac{1}{12}$, $c_{1}=\frac{c_{1}^{0}}{u_{0}^{2}}$, $u_{0}=\langle\frac{\mathrm{Re}\mathrm{Tr}\sum_{x,\mu<\nu}\mathcal{P}^{U}_{\mu\nu}(x)}{6N_{c}\tilde{V}}\rangle^{1/4}$ is the tadpole improvement factor.
The TITLC fermion action uses 1-step stout smeared link $V$ with smearing parameter $\rho=0.125$,
$S_{q}(m)=\sum_{x,\mu=1,...,4,\eta=\pm}\bar{\psi}(x)\sum\frac{1+\eta\gamma_{\mu}}{2}V_{\eta\mu}(x)\psi(x{+\eta\hat{\mu}a})$
$\quad+\sum_{x}\psi(x)\left[-(4+ma)\delta_{y,x}+c_{\rm sw}\sigma^{\mu\nu}g_{0}F_{\mu\nu}\right]\psi(x),$ (18)
where $c_{\rm sw}=\frac{1}{v^{3}_{0}}$ with $v_{0}=\langle\frac{\mathrm{Re}\mathrm{Tr}\sum_{x,\mu<\nu}\mathcal{P}^{V}_{\mu\nu}(x)}{6N_{c}\tilde{V}}\rangle^{1/4}$, and $F_{\mu\nu}$ has the same form as (21). The parameters used for the simulation are described in Table 1.
.2 Bare nucleon matrix elements with topological current in the moving frame
In lattice QCD computations of the topological current,
$K^{\mu}(x)=\epsilon^{\mu\nu\rho\sigma}\text{Tr}[A_{\nu}F_{\rho\sigma}-\frac{2}{3}ig_{0}A_{\nu}A_{\rho}A_{\sigma}](x),$ (19)
the fundamental components requiring evaluation are the gauge potential $A_{\mu}$ and the field strength tensor $F_{\mu\nu}$, defined through the following discretized formulations:
$A_{\mu}(x+\hat{\mu}/2)=\left[\frac{U_{\mu}(x)-U_{\mu}^{\dagger}(x)}{2ig_{0}a}\right]_{\mathrm{traceless}.},$ (20)
$F_{\mu\nu}$ $=$ $\frac{i}{8g_{0}a^{2}}(\mathcal{P}_{\mu,\nu}-\mathcal{P}_{\nu,\mu}+\mathcal{P}_{\nu,-\mu}-\mathcal{P}_{-\mu,\nu}$ (21)
$+\mathcal{P}_{-\mu,-\nu}-\mathcal{P}_{-\nu,-\mu}+\mathcal{P}_{-\nu,\mu}-\mathcal{P}_{\mu,-\nu}),$
where $U_{\mu}(x)$ denotes the gauge link variable and $\mathcal{P}_{\mu,\nu}=U_{\mu}(x)U_{\nu}(x+a\hat{\mu})U^{\dagger}_{\mu}(x+a\hat{\nu})U^{\dagger}_{\nu}(x)$ represents the plaquette operator. An alternative formulation employs the full-lattice definition of the gauge potential, $\tilde{A}_{\mu}(x)=(A_{\mu}(x+\hat{\mu}/2)+A_{\mu}(x-\hat{\mu}/2))/2$. The distinction between half-lattice and full-lattice definitions manifests primarily as different discretization errors. Consequently, we adopt the half-lattice definition of the gauge potential for all subsequent calculations.
Based on the definition of local topological current in (19), one can give the expression of $K^{\mu}$ in different components, namely
$K^{i}$ $\equiv$ $K^{i}_{E\times A}+K^{i}_{A_{t}B}+K^{i}_{3A}=2\text{Tr}[A\times E]+\epsilon^{tjk}\text{Tr}[A_{t}B_{k}]-\frac{2}{3}ig_{0}\epsilon^{jkt}\mathrm{Tr}[A_{j}A_{k}A_{t}],$
$K^{t}$ $\equiv$ $K^{t}_{B\cdot A}+K^{t}_{3A}=\text{Tr}[A\cdot B]-\frac{2}{3}ig_{0}\epsilon^{ijk}\mathrm{Tr}[A_{i}A_{j}A_{k}],$ (22)
where $E_{i}=F_{ti}$ and $B_{i}=F_{jk}$. The temporal component $K^{t}$, being predominantly determined by the magnetic field strength term $K^{t}_{B\cdot A}$, exhibits matrix elements that are consistent with zero within statistical uncertainties at zero momentum, yet demonstrate a gradual increase followed by saturation to a constant value with rising momentum. Conversely, the spatial components $K^{i}$, which are primarily governed by the electric field strength term $K^{i}_{E\times A}$, maintain approximately constant matrix elements within error margins across different momentum values. The physical interpretation for this behavior aligns with (16), stemming from the treatment of protons as point-like entities: the charge density represents an intrinsic property independent of momentum, while the current density evolves from zero at rest toward a limiting value with increasing momentum. These characteristic momentum dependencies constitute a direct manifestation of Lorentz covariance within a specific reference frame.
To establish the presence of Lorentz covariance, we compute the matrix elements of the topological current operator for proton external states with varying momenta. This investigation requires two complementary computational stages: first, obtaining a stable plateau in the effective mass extraction from proton two-point correlation functions; second, calculating the disconnected three-point correlation functions with topological current insertions. Within the lattice QCD framework, appropriate smearing techniques must be employed to achieve adequate signal-to-noise ratios for both proton mass determinations and $K^{\mu}$ matrix element extractions.
FIG. 4 presents the proton effective mass behavior under ensemble C24P29. Analysis reveals that while the distillation smearing technique alone provides well-optimized plateau regions for the effective mass at low momenta, this stability deteriorates significantly at higher momenta (e.g., p=1.96GeVp=1.96\ $\mathrm${GeV}). In comparison, implementation of distillation + momentum smearing scheme yields moderately improved plateau stability at large momenta relative to the distillation-only approach.
Refer to caption Refer to caption
Figure 4: Proton effective mass with different momentum under different smearing schemes under ensemble C24P29. Left figure represents the distillation smearing scheme alone, and the right figure represents the distillation + momentum smearing. The red, green and blue points represent p=0,0.98,1.96GeVp=0,0.98,1.96\ $\mathrm${GeV}, respectively.
Employing the identical smearing methodology as proton effective mass, we extract the BMEs of $K^{t/i}$ under ensemble C24P29. Ratio data presented in FIG. 5 serve to validate the theoretical discussion surrounding (.2). In summary, the employed smearing technique demonstrates robust performance throughout the computational methodology.
Refer to caption Refer to caption
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 5: Ratio $R_{K^{\mu}}(t_{f},t_{i})$ at $t_{f}\in[0.63,1.16]\ \mathrm{fm}$ and the BMEs of $K^{\mu}/(2E)$ (the gray error band) for proton external states with varying momenta under ensemble C24P29, utilizing the fitting procedure described in (9). The results are organized such that the first column corresponds to spatial components $K^{i}$, while the second column represents temporal component $K^{t}$. Rows one through three display results for proton momenta of p=0,0.98,1.96GeVp=0,0.98,1.96\ $\mathrm${GeV}, respectively.
.3 Non-perturbative renormalization with cluster decomposition error reduction (CEDR)
The computation of the renormalization constant $Z_{11}^{\mathrm{RI}}$ within the $\mathrm{RI/MOM}$ scheme, as defined in (14), necessitates the evaluation of the three-point correlation function specified in (10). This calculation presents significant numerical challenges, even with precise determinations of the gluon propagator (11) at better than $1\%$ accuracy level. The extent of this difficulty is demonstrated in FIG. 2, where the light background bands represent direct computations of $Z_{11}^{\mathrm{RI}}$ based on approximately 400 configurations of the $n_{f}=2+1$ TITLC + TITLS gauge ensemble C48P23 with lattice spacing $a\approx 0.1$ fm Hu:2023jet. These direct calculations exhibit substantial statistical uncertainties, rendering $Z_{11}^{\mathrm{RI}}$ effectively unresolved across all scales.
Applying the cluster decomposition error reduction (CDER) technique enables significant error suppression in $Z_{11}^{\mathrm{RI}}$ calculations Liu:2017man; Yang:2018bft. The fundamental principle of cluster decomposition asserts that correlators exhibit exponential decay with increasing separation between operator insertions, implying that integration beyond the correlation length primarily contributes to statistical noise rather than meaningful signal. The CDER methodology addresses this by truncating the volume integral beyond a characteristic distance, thereby achieving an enhancement of the signal-to-noise ratio by a factor of $\sqrt{V}$, as demonstrated in Liu:2017man. When implementing CDER for the computation in (10), we introduce two distinct cutoff parameters: $r_{1}$ between the gluon operator and one gauge field, and $r_{2}$ between the gauge fields within the gluon propagator. This implementation yields the modified correlator with appropriate spatial cutoffs. The CDER technique achieves computational efficiency of $\mathcal{O}(V\log V)$ through the repeated application of fast Fourier transforms (FFTs). For the three-point function specified in (10), we implement the following computational strategy:
1) Construct the spatially-restricted topological current operator: We compute $K^{\rho}_{r^{\prime}}(x)=\int_{|r^{\prime}|<r_{2}}d^{4}r^{\prime}K^{\rho}(x+r^{\prime})$ by first performing Fourier transforms (FT) of both $K^{\rho}(x)$ and the Heaviside function $f(x)=\theta(r_{2}-|x|)$, multiplying the transformed functions in momentum space, and subsequently applying anti-FT.
2) Form the intermediate operator product: We calculate $B^{\rho}_{\mu,r^{\prime}}(x)=A_{\mu}(x)K^{\rho}_{r^{\prime}}(x)$ through direct multiplication in coordinate space.
3) Apply cluster decomposition to the correlation function: For the expression $\int d^{4}xd^{4}ye^{ip\cdot(x-y)}B^{\rho}_{\mu,r^{\prime}}(x)A_{\nu}(y)$, we perform FT of both $A_{\nu}$ and $B^{\rho}_{\mu,r^{\prime}}$, apply anti-FT to the product $A_{\nu}(p)B^{\rho}_{\mu,r^{\prime}}(-p)$, implement an additional Heaviside function $g(x)=\theta(r_{1}-|x|)$ in coordinate space, and FT the resulting product.
4) Get a cutoff bare Green’s function in this way:
$G_{K}^{\mu\rho\nu,\mathrm{CDER}}(p,p)=\langle\int_{|r|<r_{1}}d^{4}r\int_{|r^{\prime}|<r_{2}}d^{4}r^{\prime}\int d^{4}xe^{ip\cdot r^{\prime}}K^{\rho}(x+r^{\prime})\mathrm{Tr}[A_{\mu}(x)A_{\nu}(x+r)]\rangle.$ (23)
Here both momentum arguments of bare Green’s function are set to $p$ in accordance with the standard implementation of the $\mathrm{RI/MOM}$ scheme, where the renormalization scale is defined by $\mu_{\rm RI}^{2}=p^{2}$.
The selection of cutoff parameters $r_{1,2}$ in the CDER technique requires careful consideration, as inappropriate choices may induce periodic oscillations in the renormalization constants Liu:2017man. FIG. 6 provides visual confirmation of this analysis. When the cutoff parameter governing the gluon field-operator correlation ($r_{1}$) is set too small (top left panel), distinct periodic oscillations become apparent in the results. These oscillations progressively stabilize as $r_{1}$ increases. A similar pattern is observed for variations in the current cutoff parameter ($r_{2}$), though changes in oscillatory behavior is less pronounced without corresponding adjustments to $r_{1}$. In FIG. 2, we employ the optimal cutoffs identified in the bottom right panel of FIG. 6, specifically $r_{1}=1.26\ \mathrm{fm}$ and $r_{2}=0.74\ \mathrm{fm}$, to perform discretization error subtraction procedure. Thus, we have successfully computed the purely gluonic Green’s function using the CDER technique, which is crucial for the non-perturbative renormalization of topological current. This method effectively suppresses the significant statistical noise generated when directly computing the three-point correlation function by truncating the volume integral beyond the characteristic distances $r_{1}$ and $r_{2}$, thus enabling the stable determination of the renormalization constant $Z_{11}^{\mathrm{RI}}$, which was previously only obtainable through the stacking statistics.
Refer to caption Refer to caption
Refer to caption Refer to caption
Figure 6: The $\overline{\mathrm{MS}}$ renormalization constant $\bar{Z}_{11}^{\mathrm{lat}}(a^{2}p^{2},{\mu}^{2}=10\ \mathrm{GeV}^{2})$ under ensemble C48P23 as a function of $a^{2}p^{2}$ using momentum modes $(p_{\mathrm{lat}},p_{\mathrm{lat}},p_{\mathrm{lat}},p_{\mathrm{lat}})$, with different cutoffs on the gluon field-operator correlation ($r_{1}$) and current ($r_{2}$).
.4 Three-loop approximation of $\Delta G$’s off-diagonal elements
This subsection outlines the derivation of the reduced expressions presented in (15). Starting from (Total Gluon Helicity Contribution to Proton Spin from Lattice QCD) and (Total Gluon Helicity Contribution to Proton Spin from Lattice QCD), we obtain the exact formulations for gluon helicity in the $\overline{\mathrm{MS}}$ scheme:
$\Delta G^{\overline{\mathrm{MS}}}(a^{2}p^{2},{\mu}^{2})$ $=$ $R_{11}(a^{2}p^{2},{\mu}^{2})Z_{11}^{\mathrm{RI}}(a^{2}p^{2})\big(\Delta G^{\mathrm{B.}}-\frac{K^{\mathrm{lat.,q}}}{J_{5}^{\mathrm{lat.,q}}}(a^{2}p^{2})\Delta\Sigma^{\mathrm{B.}}\big)$ (24)
$+R_{12}(a^{2}p^{2},{\mu}^{2})Z_{A}^{\mathrm{RI}}(a^{2}p^{2})\Delta\Sigma^{\mathrm{B.}}.$
In this formulation, we express the dependence on the squared renormalization scale $\mu_{\rm RI}^{2}$ of the $\mathrm{RI/MOM}$ scheme in terms of the dimensionless quantity $a^{2}p^{2}$, and the same below. We now implement a three-loop approximation for the ratio term $K^{\mathrm{lat,q}}/{J_{5}^{\mathrm{lat,q}}}$. Noting that $R_{12}$ can be determined from the topological current $K^{\mu}$ under quark external states while $R_{22}\approx 1$ is fixed by the axial vector current under quark external states, we employ the approximation:
$\big(\bar{Z}_{11}^{\mathrm{lat}}R_{12}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2})$ $\approx$ $\big(\bar{Z}_{11}^{\mathrm{lat}}\frac{R_{12}}{R_{22}}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2})$ (25)
$=$ $\big(\bar{Z}_{11}^{\mathrm{lat}}\frac{K^{3\mathrm{-loop,q}}}{J_{5}^{3\mathrm{-loop,q}}}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2})$
$\approx$ $\big(\bar{Z}_{11}^{\mathrm{lat}}\frac{K^{\mathrm{lat,q}}}{J_{5}^{\mathrm{lat,q}}}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2}).$
where we used a redefined expression $\bar{Z}_{11}^{\mathrm{lat}}\equiv R_{11}Z_{11}^{\mathrm{RI}}$. Furthermore, owing to the protection by chiral symmetry, the matching coefficient for the axial-vector current remains unity up to two loops, i.e., $\bar{Z}_{A}^{\mathrm{lat}}\equiv Z_{A}^{\overline{\mathrm{MS}}}\approx Z_{A}^{\mathrm{RI}}$. This convention will be consistently maintained throughout the subsequent discussion. This approximation enables the following simplified expression for gluon helicity at three-loop order:
$\Delta G^{\overline{\mathrm{MS}}}$ $\overset{3\mathrm{-loop}}{\approx}$ $\bar{Z}_{11}^{\mathrm{lat}}\big(\Delta G^{\mathrm{B.}}-R_{12}\Delta\Sigma^{\mathrm{B}.}\big)+R_{12}\bar{Z}_{A}^{\mathrm{lat}}\Delta\Sigma^{\mathrm{B.}}$ (26)
$=$ $\bar{Z}_{11}^{\mathrm{lat}}\Delta G^{\mathrm{B.}}+(\bar{Z}_{A}^{\mathrm{lat}}-\bar{Z}_{11}^{\mathrm{lat}})R_{12}\Delta\Sigma^{\mathrm{B.}}$
$\approx$ $\bar{Z}_{11}^{\mathrm{lat}}\Delta G^{\mathrm{B.}},$
where the final reduction is justified based on the hierarchy $\bar{Z}_{11}^{\mathrm{lat}}\gg(\bar{Z}_{A}^{\mathrm{lat}}-\bar{Z}_{11}^{\mathrm{lat}})R_{12}$, and by the fact that $\Delta G^{\mathrm{B.}}$ is comparable in magnitude to $\Delta\Sigma^{\mathrm{B.}}$. The derivation of expression $\Delta\Sigma^{\overline{\mathrm{MS}}}\approx\bar{Z}_{A}^{\mathrm{lat}}\Delta\Sigma^{\mathrm{B.}}$ parallels the approach used in the approximation above.
.5 Perturbative results for matching coefficients $R_{ij}$ up to three-loops
This section details the perturbative calculation of the finite renormalization (or matching) coefficients $R_{ij}$ introduced in eq. (Total Gluon Helicity Contribution to Proton Spin from Lattice QCD) for the singlet axial-vector current $J_{5}^{\mu}$ and topological current $K^{\mu}$ from the intermediate $\mathrm{RI/MOM}$ scheme used in Lattice calculation to the $\overline{\mathrm{MS}}$ scheme typically employed in perturbative calculations in Dimensional Regularization (DR). The explicit results for $R_{ij}$ up to three loops are documented for the sake of references.
In the spirit of $\mathrm{RI/MOM}$, the finite matching coefficients $R_{ij}$ can be defined in terms of the operator renormalization constants as follows:
$\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}\equiv\begin{pmatrix}\bar{Z}^{\mathrm{DR}}_{11}&\bar{Z}^{\mathrm{DR}}_{12}\\ \bar{Z}^{\mathrm{DR}}_{21}&\bar{Z}^{\mathrm{DR}}_{22}\end{pmatrix}\cdot\begin{pmatrix}Z^{\mathrm{RI,\,DR}}_{11}&Z^{\mathrm{RI,\,DR}}_{12}\\ Z^{\mathrm{RI,\,DR}}_{21}&Z^{\mathrm{RI,\,DR}}_{22}\end{pmatrix}^{-1}\Big|_{\epsilon\rightarrow 0}=\begin{pmatrix}\bar{Z}^{\mathrm{lat}}_{11}&\bar{Z}^{\mathrm{lat}}_{12}\\ \bar{Z}^{\mathrm{lat}}_{21}&\bar{Z}^{\mathrm{lat}}_{22}\end{pmatrix}\cdot\begin{pmatrix}Z^{\mathrm{RI,\,lat}}_{11}&Z^{\mathrm{RI,\,lat}}_{12}\\ Z^{\mathrm{RI,\,lat}}_{21}&Z^{\mathrm{RI,\,lat}}_{22}\end{pmatrix}^{-1}\,.$ (27)
Consequently, the $R_{ij}$ determined using the first equality $\bar{Z}^{\mathrm{DR}}\cdot\big(Z^{\mathrm{RI,\,DR}}\big)^{-1}\big|_{\epsilon\rightarrow 0}$ with DR can be employed, together with the input $Z^{\mathrm{RI,\,lat}}_{ij}\equiv Z^{\mathrm{RI}}_{ij}$ computed in (Total Gluon Helicity Contribution to Proton Spin from Lattice QCD), to obtain the values for the $\overline{\mathrm{MS}}$-renormalization constants $\bar{Z}^{\mathrm{lat}}_{ij}$ of the operator $K^{\mu}$ and $J_{5}^{\mu}$ but with UV-divergences regularized using Lattice, which can then be applied directly to the bare operator matrix elements computed in Lattice to obtain the $\overline{\mathrm{MS}}$-renormalized results. By virtue of (3), we can further derive the following shortcut expressions for $R_{ij}$:
$\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}=\begin{pmatrix}\frac{\overline{K}_{gg}}{K^{\mathrm{RI}}_{gg}}&\frac{\overline{K}_{\bar{q}q}}{J^{\mathrm{RI}}_{\bar{q}q}}\\ \frac{\overline{J}_{gg}}{K^{\mathrm{RI}}_{gg}}&\frac{\overline{J}_{\bar{q}q}}{J^{\mathrm{RI}}_{\bar{q}q}}\end{pmatrix}$ (28)
where $\overline{K}(\overline{J})_{gg(\bar{q}q)}$ denote the Lorentz-invariant light-cone form-factors or matrix elements of the $\overline{\mathrm{MS}}$-renormalized $K(J)$-current operator but with $\mathrm{RI/MOM}$-renormalized external $g(q)$-fields. Note that $K^{\mathrm{RI}}_{gg}$ and $J^{\mathrm{RI}}_{\bar{q}q}$ equal to, respectively, the tree-level values of the aforementioned matrix elements of $K(J)$-current operator as a result of the $\mathrm{RI/MOM}$ renormalization condition (3).
Extra care is required in the calculation of the dimensionally-regularized bare matrix elements and the subsequent extraction of renormalization constants entering the r.h.s. of (27) and (28), because of the notorious issue in treating $\gamma_{5}$, and to less extent $\epsilon^{\mu\nu\rho\sigma}$ in DR, especially so in the present case due to the involvement of axial anomaly. In view of the applications related to the study of polarized structure functions of the proton, we choose a (modified) $\overline{\mathrm{MS}}$ prescription where the gauge invariance and Adler-Bell-Jackiw equation are preserved (hence without chiral symmetry). So far, there is no practical prescription known to the authors to apply consistently an anticommuting $\gamma_{5}$ in DR for computing loop amplitudes with (closed) fermion chains with an odd number of $\gamma_{5}$, especially when the amplitudes are anomalous Chen:2023lus; Chen:2024hlv. On the other hand, the so-called Larin’s prescription Larin:1993tq of $\gamma_{5}$, where $\gamma_{5}$ in DR is treated non-anticommuting tHooft:1972tcz; Breitenlohner:1977hr; Breitenlohner:1975hg; Breitenlohner:1976te while the spacetime-metric tensors resulting from contracting a pair of $\epsilon^{\mu\nu\rho\sigma}$ are all in D-dimensions, is particularly suited for dealing with the QCD corrections at hand: both gauge invariance and the Adler-Bell-Jackiw equation are maintained and it is technically convenient to apply at high loop orders in QCD.
With this set-up, we have derived the perturbative expressions for the bare matrix elements involved in (28) up to three loops, truncated to high orders in $\epsilon$ sufficient for obtaining their renormalized counterparts to $\mathcal{O}(\epsilon^{0})$; they are subsequently employed to extract the perturbative results for the finite $R_{ij}$ according to both (28) and (27), with perfect agreement found between the two. (For more technical details of the computational workflow and evaluation of loop integrals, we kindly refer to refs. Ahmed:2021spj; Chen:2021gxv.) For the sake of references, we document below the perturbative results for the finite matching coefficients $R_{ij}$ up to three loops in Landau gauge.
We begin with $R_{11}$ that was utilized in the present work, which reads
$R_{11}$ $=$ $1\,+\,a_{s}\,\Big(C_{A}\,\left(\frac{11L_{\mu}}{3}+\frac{367}{36}\right)\,+\,n_{f}\,\left(-\frac{2L_{\mu}}{3}-\frac{10}{9}\right)\Big)$ (29)
$+$ $a_{s}^{2}\,\Big(C_{A}^{2}\,\left(\frac{15\zeta_{3}}{4}+\frac{121L_{\mu}^{2}}{9}+\frac{4649L_{\mu}}{54}+\frac{99047}{648}\right)\,+\,C_{A}n_{f}\,\left(-16\zeta_{3}-\frac{44L_{\mu}^{2}}{9}-\frac{677L_{\mu}}{27}-\frac{2762}{81}\right)$
$+$ $C_{F}n_{f}\,\left(16\zeta_{3}-8L_{\mu}-\frac{161}{6}\right)\,+\,n_{f}^{2}\,\left(\frac{4L_{\mu}^{2}}{9}+\frac{40L_{\mu}}{27}+\frac{100}{81}\right)\Big)$
$+$ $a_{s}^{3}\,\Big(C_{A}^{3}\,\left(-\frac{3743\zeta_{3}}{288}-\frac{31475\zeta_{5}}{192}+\frac{165\zeta_{3}L_{\mu}}{4}+\frac{1331L_{\mu}^{3}}{27}+\frac{55627L_{\mu}^{2}}{108}+\frac{1273537L_{\mu}}{648}+\frac{17627903}{5832}\right)$
$+$ $C_{A}^{2}n_{f}\,\left(-\frac{16627\zeta_{3}}{36}+\frac{470\zeta_{5}}{3}-\frac{367\zeta_{3}L_{\mu}}{2}-\frac{242L_{\mu}^{3}}{9}-\frac{2194L_{\mu}^{2}}{9}-\frac{28805L_{\mu}}{36}-\frac{2061755}{1944}\right)$
$+$ $C_{A}C_{F}n_{f}\,\left(\frac{1306\zeta_{3}}{3}+120\zeta_{5}+176\zeta_{3}L_{\mu}-\frac{187L_{\mu}^{2}}{3}-\frac{2735L_{\mu}}{6}-\frac{331903}{324}\right)$
$+$ $C_{A}n_{f}^{2}\,\left(\frac{524\zeta_{3}}{9}+32\zeta_{3}L_{\mu}+\frac{44L_{\mu}^{3}}{9}+\frac{319L_{\mu}^{2}}{9}+\frac{4895L_{\mu}}{54}+\frac{89833}{972}\right)$
$+$ $C_{F}^{2}n_{f}\,\left(\frac{460\zeta_{3}}{3}-240\zeta_{5}+19L_{\mu}+\frac{1333}{18}\right)\,+\,C_{F}n_{f}^{2}\,\left(-72\zeta_{3}-32\zeta_{3}L_{\mu}+\frac{34L_{\mu}^{2}}{3}+\frac{202L_{\mu}}{3}+\frac{20375}{162}\right)$
$+$ $n_{f}^{3}\,\left(-\frac{8L_{\mu}^{3}}{27}-\frac{40L_{\mu}^{2}}{27}-\frac{200L_{\mu}}{81}-\frac{1000}{729}\right)\Big)\,+\,\mathcal{O}(a_{s}^{4})\,$,
where the shorthand notations $a_{s}\equiv\frac{\alpha_{s}}{4\,\pi}$ and $L_{\mu}\equiv\ln\big(\frac{{\mu}^{2}}{\mu_{\rm RI}^{2}}\big)$ are introduced. The definitions of the quadratic Casimir color constants involved above are as usual: $C_{A}=N_{c}\,,\,C_{F}=(N_{c}^{2}-1)/(2N_{c})\,$ with $N_{c}=3$ in QCD and we have set the color-trace normalization factor to its value 1/2{1}/{2}. ($\zeta_{n}$ is the usual Riemann $\zeta(n)$ function.) The perturbative corrections to $R_{22}$ starts only from two-loops and are simpler:
$R_{22}$ $=$ $1\,+\,a_{s}^{2}\,\Big(C_{F}n_{f}\,\left(-6L_{\mu}-12\right)\Big)$ (30)
$+$ $a_{s}^{3}\,\Big(C_{A}C_{F}n_{f}\,\left(66\zeta_{3}-22L_{\mu}^{2}-\frac{406L_{\mu}}{3}-\frac{4643}{18}\right)$
$+$ $C_{F}^{2}n_{f}\,\left(-48\zeta_{3}+18L_{\mu}+\frac{77}{2}\right)\,+\,C_{F}n_{f}^{2}\,\left(4L_{\mu}^{2}+\frac{52L_{\mu}}{3}+\frac{289}{9}\right)\Big)\,+\,\mathcal{O}(a_{s}^{4})\,.$
The results for the off-diagonal elements are:
$R_{12}$ $=$ $1\,+\,a_{s}\,\Big(C_{F}\,\left(3L_{\mu}+6\right)\Big)$ (31)
$+$ $a_{s}^{2}\,\Big(C_{A}C_{F}\,\left(-33\zeta_{3}+11L_{\mu}^{2}+\frac{203L_{\mu}}{3}+\frac{4643}{36}\right)\,+\,C_{F}^{2}\,\left(24\zeta_{3}-9L_{\mu}-\frac{77}{4}\right)\,+\,C_{F}n_{f}\,\left(-2L_{\mu}^{2}-\frac{26L_{\mu}}{3}-\frac{289}{18}\right)\Big)$
$+$ $a_{s}^{3}\,\Big(C_{A}^{2}C_{F}\,\left(-\frac{14419\zeta_{3}}{12}+\frac{1045\zeta_{5}}{4}-363\zeta_{3}L_{\mu}+\frac{121L_{\mu}^{3}}{3}+\frac{2437L_{\mu}^{2}}{6}+\frac{30395L_{\mu}}{18}+\frac{1691641}{648}\right)$
$+$ $C_{A}C_{F}^{2}\,\left(668\zeta_{3}+264\zeta_{3}L_{\mu}-\frac{99L_{\mu}^{2}}{2}-327L_{\mu}-\frac{104701}{216}\right)$
$+$ $C_{A}C_{F}n_{f}\,\left(\frac{140\zeta_{3}}{3}-6\zeta_{3}L_{\mu}-\frac{44L_{\mu}^{3}}{3}-\frac{376L_{\mu}^{2}}{3}-\frac{4148L_{\mu}}{9}-\frac{2\pi^{4}}{5}-\frac{102293}{162}\right)$
$+$ $C_{F}^{3}\,\left(64\zeta_{3}-240\zeta_{5}+\frac{63L_{\mu}}{2}+\frac{707}{12}\right)\,+\,C_{F}^{2}n_{f}\,\left(72\zeta_{3}+24\zeta_{3}L_{\mu}-6L_{\mu}^{2}-75L_{\mu}+\frac{2\pi^{4}}{5}-\frac{12899}{54}\right)$
$+$ $C_{F}n_{f}^{2}\,\left(\frac{8\zeta_{3}}{3}+\frac{4L_{\mu}^{3}}{3}+\frac{26L_{\mu}^{2}}{3}+\frac{250L_{\mu}}{9}+\frac{5005}{162}\right)\Big)\,+\,\mathcal{O}(a_{s}^{4})\,$,
and
$R_{21}$ $=$ $1\,+\,a_{s}\,n_{f}\,\left(-2\right)\,+\,a_{s}^{2}\,\Big(C_{A}n_{f}\,\left(-\frac{22L_{\mu}}{3}-\frac{367}{18}\right)\,+\,n_{f}^{2}\,\left(\frac{4L_{\mu}}{3}+\frac{20}{9}\right)\Big)$ (32)
$+$ $a_{s}^{3}\,\Big(C_{A}^{2}n_{f}\,\left(-\frac{15\zeta_{3}}{2}-\frac{242L_{\mu}^{2}}{9}-\frac{4649L_{\mu}}{27}-\frac{99047}{324}\right)\,+\,C_{A}n_{f}^{2}\,\left(32\zeta_{3}+\frac{88L_{\mu}^{2}}{9}+\frac{1354L_{\mu}}{27}+\frac{5524}{81}\right)$
$+$ $C_{F}n_{f}^{2}\,\left(-32\zeta_{3}+16L_{\mu}+\frac{161}{3}\right)\,+\,n_{f}^{3}\,\left(-\frac{8L_{\mu}^{2}}{9}-\frac{80L_{\mu}}{27}-\frac{200}{81}\right)\Big)\,+\,\mathcal{O}(a_{s}^{4})\,.$
The dependence of $R_{ij}$ on ${\mu}^{2}$ exhibited in the above perturbative expressions are governed by their RG equations. Following from the definition (27) the RG equations for the matching coefficients $R_{ij}$ read
${\mu}^{2}\,\frac{\mathrm{d}}{\mathrm{d}\,{\mu}^{2}}\,\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}=\begin{pmatrix}\gamma_{11}&\gamma_{12}\\ 0&\gamma_{22}\end{pmatrix}\cdot\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}\,$, (33)
where the matrix of anomalous dimensions in r.h.s. is fully determined by the (proper or modified) $\overline{\mathrm{MS}}$ renormalization constants in (27) in Larin’s prescription Larin:1993tq; these anomalous dimensions observe the relations Breitenlohner:1983pi; Ahmed:2021spj; Luscher:2021bog $\gamma_{11}=-\frac{\mathrm{d}\ln a_{s}}{\mathrm{d}\ln{\mu}^{2}}=-\beta$ and $\gamma_{22}=-2a_{s}\,n_{f}\,\gamma_{12}$ known now to four Chen:2021gxv and five loops Chen:2022lun
With the explicit expression (29) for $R_{11}$, it is interesting to observe that in the limit of large $\beta_{0}$ or $n_{f}$, the leading terms in powers of $\mathcal{O}(a^{n}_{s}\,\beta^{n}_{0})$ exhibit the following neat pattern:
$R_{11}\Big|_{\text{large $\beta_{0}$}}\approx 1\,+\,a_{s}\,\beta_{0}\,\big(L_{\mu}+\frac{5}{3}\big)\,+\,a_{s}^{2}\,\beta_{0}^{2}\,\big(L_{\mu}+\frac{5}{3}\big)^{2}\,+\,a_{s}^{3}\,\beta_{0}^{3}\,\big(L_{\mu}+\frac{5}{3}\big)^{3}\,+\,\mathcal{O}(a_{s}^{4})\,$, (34)
[…]
Refer to caption Refer to caption
Figure 7: Three-loop (left panel) and two-loop (right panel) matching coefficients in the $\mathrm{RI/MOM}$ scale at ${\mu}^{2}=10\ \mathrm{GeV}^{2}$. In order to combine with lattice data, we change the horizontal axis to the dimensionless $\mathrm{RI/MOM}$ scale $a^{2}p^{2}$.
Figure 8: One to three-loop matching coefficients $R_{11}$ in the $\mathrm{RI/MOM}$ scale at ${\mu}^{2}=10\ \mathrm{GeV}^{2}$. In order to combine with lattice data, we change the horizontal axis to the dimensionless $\mathrm{RI/MOM}$ scale $a^{2}p^{2}$.
Table 2: Raw data employed for the global fitting corresponding to (16). The total uncertainty in the renormalization constants (for both the two- and three-loop matching) comprises two components: the statistical error and systematic error, which the systematic error given by difference between the center value of RG-evolved result and fixed-order result.
$a$ (fm) Bare matrix elements $\bar{Z}_{11}^{\mathrm{lat}}$
$p(\mathrm{GeV})$ $\langle K^{t}\rangle/(2E)$ $\langle K^{i}\rangle/(2E)$ $2\mathrm{-loop}$ $3\mathrm{-loop}$
0.10524(05)(62) 0.00 0.006(05) 0.212(30) 0.91(09)(03) 1.09(12)(04)
0.49 0.110(18) 0.209(11)
0.98 0.154(25) 0.187(26)
1.47 0.195(38) 0.180(27)
1.96 0.212(40) 0.202(40)
2.45 0.214(46) 0.152(93)
0.08973(20)(53) 0.00 0.004(07) 0.164(07) 0.89(10)(04) 1.07(12)(04)
0.86 0.193(16) 0.202(53)
1.29 0.214(53) 0.171(63)
0.07753(03)(45) 0.00 0.001(07) 0.166(09) 0.85(05)(03) 0.98(06)(03)
1.00 0.181(10) 0.165(12)
1.50 0.173(30) 0.176(32)

### v2

@@
-Total Gluon Helicity Contribution to Proton Spin from Lattice QCD
+Total Gluon Helicity Contribution to the Proton Spin from Lattice QCD
 Abstract
-We report a state-of-the-art lattice QCD calculation of the total gluon helicity contribution to proton spin, $\Delta G$.
-The calculation is done on ensembles at three different lattice spacings $a=\{0.08,0.09,0.11\}$ fm.
-By employing distillation + momentum smearing for proton external states, we extract the bare matrix elements of the topological current $K^{\mu}$ under the 5-HYP smeared Coulomb gauge fixing configurations.
-Furthermore, we apply a non-perturbative $\mathrm{RI/MOM}$ renormalization scheme augmented with the Cluster Decomposition Error Reduction (CDER) technique to determine the renormalization constants of $K^{\mu}$.
+We report a state-of-the-art lattice QCD calculation of the total gluon helicity contribution to the proton spin, $\Delta G$.
+The calculation is done on ensembles with three different lattice spacings $a=\{0.08,0.09,0.11\}$ fm.
+By employing distillation and momentum smearing for proton external states, we extract the bare matrix elements of the topological current $K^{\mu}$ using 5-HYP smeared Coulomb gauge fixing configurations.
+Furthermore, we apply a non-perturbative $\mathrm{RI/MOM}$ renormalization scheme augmented by the Cluster Decomposition Error Reduction (CDER) technique to determine the renormalization constants of $K^{\mu}$.
 The results obtained from different components $K^{t,i}$ (with $i$ being the direction of proton momentum or polarization) are consistent with Lorentz covariance within uncertainties.
 After extrapolating to the continuum limit, $\Delta G$ is found to be $\Delta G=0.231(17)^{\mathrm{sta.}}(44)^{\mathrm{sym.}}$ at the $\overline{\mathrm{MS}}$ scale ${\mu}^{2}=10\ \mathrm{GeV}^{2}$, which constitutes approximately $46(9)\%$ of the proton spin.
-Introduction: Since the European Muon Collaboration (EMC) measured the quark spin contribution to proton spin through deep inelastic scattering with polarized muon and proton 1988364, the origin of proton spin has remained one of the most profound puzzles in modern particle and nuclear physics Ji:2020ena.
-Our current understanding is that the quark spin contribution accounts for only $\sim 30\%$ of the proton spin deFlorian:2009vb; Nocera:2014gqa; COMPASS:2015mhb; Alexandrou:2017oeh; Lin:2018obj; Liang:2018pis.
+Introduction: Since the European Muon Collaboration (EMC) measured the quark spin contribution to the proton spin using deep inelastic scattering of polarized muons and protons et al (1988), the origin of proton spin has remained one of the most profound puzzles in modern particle and nuclear physics Ji et al.
+(2021b).
+Our current understanding is that the quark spin contribution accounts for only $\sim 30\%$ of the proton spin de Florian et al.
+(2009); Nocera et al.
+(2014); Adolph and others (2016); Alexandrou et al.
+(2017); Lin et al.
+(2018); Liang et al.
+(2018).
 The remaining contributions come from quark orbital angular momentum (OAM), as well as gluon spin and OAM.
-Two popular spin sum rules have been proposed in the literature to understand the proton spin structure.
-One is the Jaffe-Manohar sum rule Jaffe:1989jz which offers a complete decomposition of the proton spin into quark and gluon spin and OAM.
+Two popular spin sum rules have been proposed to help understanding the proton spin structure.
+One is the Jaffe-Manohar sum rule Jaffe and Manohar (1990) which offers a complete decomposition of the proton spin into quark and gluon spin and OAM.
 It has a clear partonic picture in the infinite momentum frame (IMF) of the proton and lightcone gauge.
-The other is the frame- and gauge-independent sum rule by Ji Ji:1996ek, where the quark and gluon total angular momentum can be connected to the moments of generalized parton distributions Muller:1994ses; Ji:1996ek.
-Among these contributions, the understanding of gluon spin has posed a significant challenge, mainly due to the absence of a local gauge-invariant gluon spin operator Ji:2012gc.
-Nevertheless, it can be measured in polarized scattering experiments through the spin-dependent gluon helicity distribution $\Delta g(x)$ as its lowest moment $\Delta G=\int dx\Delta g(x)$.
+The other is the frame- and gauge-independent sum rule by Ji Ji (1997), where the quark and gluon total angular momentum can be connected to the moments of generalized parton distributions Müller et al.
+(1994); Ji (1997).
+Among these contributions, the understanding of gluon spin has posed a significant challenge, mainly due to the absence of a local gauge-invariant gluon spin operator Ji et al.
+(2012).
+Nevertheless, it can be deduced in polarized scattering experiments from the spin-dependent gluon helicity distribution $\Delta g(x)$, more precisely its lowest moment $\Delta G=\int dx\Delta g(x)$.
 Note that unlike in the case of the quark distribution, the lowest moment of the gluon helicity distribution is still nonlocal.
-Only in the lightcone gauge does it reduce to the matrix element of a local operator $\vec{E}\times\vec{A}$.
-Phenomenological global analyses of experimental data indicate that $\Delta G$ is approximately 0.2 at the $\overline{\mathrm{MS}}$ scale $\mu^{2}=10\ \mathrm{GeV}^{2}$, provided that contributions from the small-$x$ region ($x<0.05$) with large uncertainties are excluded deFlorian:2009vb; deFlorian:2014yva.
-On the other hand, recent theoretical developments Ji:2013fga; Hatta:2013gta; Ji:2014lra, particularly within the framework of Large-Momentum Effective Theory (LaMET) Ji:2013dva; Ji:2014gla; Ji:2020ect, have enabled a direct computation of $\Delta G$ using lattice QCD.
+Only in the lightcone gauge does it reduce to a matrix element of the local operator $\vec{E}\times\vec{A}$.
+Phenomenological global analyses of experimental data indicate that $\Delta G$ is approximately 0.2 at the $\overline{\mathrm{MS}}$ scale $\mu^{2}=10\ \mathrm{GeV}^{2}$, provided that contributions from the small-$x$ region ($x<0.05$) with large uncertainties are excluded de Florian et al.
+(2009, 2014).
+On the other hand, recent theoretical developments Ji et al.
+(2013); Hatta et al.
+(2014); Ji et al.
+(2015), particularly within the framework of Large-Momentum Effective Theory (LaMET) Ji (2013, 2014); Ji et al.
+(2021a), have enabled a direct computation of $\Delta G$ using lattice QCD.
 In this approach, instead of computing in the IMF and lightcone gauge which are both inaccessible on the lattice, one computes the large-momentum nucleon matrix element of a static gluon spin operator, $\vec{E}\times\vec{A}$, in a physical gauge that leaves the transverse gauge field intact.
-Actually, there exists an infinite number of operators that can be used to extract $\Delta G$, all belonging to the same universality class Hatta:2013gta.
-The static gluon spin matrix element can then be matched to $\Delta G$ through a perturbative factorization or matching relation Ji:2013fga.
-Based on this, the $\rm\chi QCD$ collaboration conducted an exploratory calculation of $\Delta G$ by evaluating the nucleon matrix element of $\vec{E}\times\vec{A}$ in the Coulomb gauge, yielding a result of $0.251(47)(16)$ Yang:2016plb.
+Actually, there exists an infinite number of operators that can be used to extract $\Delta G$, all belonging to the same universality class Hatta et al.
+(2014).
+The static gluon spin matrix element can then be matched to $\Delta G$ through a perturbative factorization or matching relation Ji et al.
+(2013).
+Based on this, the $\rm\chi QCD$ collaboration conducted an exploratory calculation of $\Delta G$ by evaluating the nucleon matrix element of $\vec{E}\times\vec{A}$ in the Coulomb gauge, which yields the result $0.251(47)(16)$ Yang et al.
+(2017).
 However, the one-loop lattice perturbative renormalization employed in that study introduces a large correction, indicating the necessity of a proper nonperturbative renormalization – which was, at the time, very challenging to implement due to the Coulomb gauge fixing.
-Furthermore, recent studies Pang:2024sdl have revealed that the LaMET factorization for $\Delta G$ is incompatible with that for $\Delta g(x)$.
+Furthermore, recent studies Pang et al.
+(2024) have revealed that the LaMET factorization for $\Delta G$ is incompatible with that for $\Delta g(x)$.
 To circumvent the difficulties mentioned above, two new proposals have been made in Ref.
-Pang:2024sdl.
+Pang et al.
+(2024).
 One is based on using the topological current, and the other employs appropriate static nonlocal gluon operators generating the gluon helicity distribution.
-Both of them have a trivial matching only, and therefore resolve the inconsistency previously encountered in the matching for $\Delta G$ and $\Delta g(x)$.
+Both of them have trivial matching only, and therefore resolve the inconsistency previously encountered in the matching for $\Delta G$ and $\Delta g(x)$.
 Moreover, their nonperturbative renormalization can be easily implemented.
 This opens a new avenue for systematic studies of $\Delta G$ from lattice QCD.
-In this work, we present a state-of-the-art lattice calculation of the total gluon helicity contribution to proton spin, $\Delta G$, using the topological current approach proposed in Ref.
-Pang:2024sdl.
+In this work, we present a state-of-the-art lattice calculation of the total gluon helicity contribution to the proton spin, $\Delta G$, using the topological current approach proposed in Ref.
+Pang et al.
+(2024).
 The calculation is done on ensembles at three different lattice spacings $a=\{0.08,0.09,0.11\}$ fm, which allow us to perform a continuum extrapolation.
-Distillation HadronSpectrum:2009krc, momentum smearing Bali:2016lva, and cluster decomposition Liu:2017man techniques have been used to improve the signal-to-noise ratio of proton matrix elements and parton Green’s functions, respectively.
+Distillation Peardon et al.
+(2009), momentum smearing Bali et al.
+(2016); Egerer et al.
+(2021), and cluster decomposition Liu et al.
+(2018) techniques have been used to improve the signal-to-noise ratio of proton matrix elements and parton Green’s functions, respectively.
 The nonperturbative renormalization has been performed in the $\mathrm{RI/MOM}$ scheme, following the operator mixing pattern given in Ref.
-Pang:2024sdl.
+Pang et al.
+(2024).
 After converting to the $\overline{\rm MS}$ scheme, the result is then evolved to the scale $\mu^{2}=10\,{\rm GeV}^{2}$ using the renormalization group (RG) equation with three-loop anomalous dimensions.
-Theoretical Framework: Following Ref. Pang:2024sdl, the total gluon helicity $\Delta G$ can be extracted by computing the matrix element of the local topological current
+Theoretical Framework: Following Ref. Pang et al. (2024), the total gluon helicity $\Delta G$ can be extracted by computing the matrix element of the local topological current
 $K^{\mu}(x)=\epsilon^{\mu\nu\rho\sigma}\text{Tr}[A_{\nu}F_{\rho\sigma}-\frac{2}{3}ig_{0}A_{\nu}A_{\rho}A_{\sigma}](x)$ (1)
-in a longitudinally polarized proton.
-More specifically, the proton matrix element of the Coulomb gauge fixed operator $K^{\mu}$ ($\mu=t,i$ with $i$ denoting the component along the direction of proton momentum or polarization) reduces to $\Delta G$ in the IMF, namely,
-$\frac{\langle\mathrm{PS}_{\mathrm{Proj.i}}|\int d^{3}xK^{t/i}_{\mathrm{C.G.}}(x)|\mathrm{PS}_{\mathrm{Proj.i}}\rangle}{2S^{t/i}}|_{\mathrm{IMF}}=\Delta G,$ (2)
-where $|\mathrm{PS}_{\mathrm{Proj.i}}\rangle$ denotes the proton external state with specific momentum and polarization direction $i$, while $S^{t/i}$ represents the corresponding polarization vector.
+in a longitudinally polarized proton. More specifically, the proton matrix element of the Coulomb gauge fixed operator $K^{\mu}$ ($\mu=t,i$) reduces to $\Delta G$ in the IMF, namely,
+$\frac{\langle\mathrm{PS}_{\mathrm{Proj.i}}|\int d^{3}xK^{\mu}_{\mathrm{C.G.}}(x)|\mathrm{PS}_{\mathrm{Proj.i}}\rangle}{2S^{\mu}\langle\mathrm{PS}_{\mathrm{Unproj.}}|\mathrm{PS}_{\mathrm{Unproj.}}\rangle}|_{\mathrm{IMF}}=\Delta G,$ (2)
+$|\mathrm{PS}_{\mathrm{Unproj.}}\rangle$ denotes the proton external state with a specific momentum but without polarization.
+In contrast, $|\mathrm{PS}_{\mathrm{Proj.i}}\rangle$ represents the state with specific momentum and a definite polarization direction $i$.
+The $\mu$ component of the corresponding polarization vector in the $i$th-direction is denoted by $S^{\mu}$.
 The subscript $\mathrm{C.G.}$ indicates the Coulomb gauge fixing for the topological current.
 As shown in Ref.
-Pang:2024sdl, the ultraviolet (UV) divergence of $K^{\mu}$ is gauge independent.
-Therefore, we can perform a nonperturbative renormalization of $K^{\mu}$ in the $\mathrm{RI/MOM}$ scheme MARTINELLI199581 where the renormalization factor is typically computed in the Landau gauge.
-$K^{\mu}$ mixes only with the axial-vector current $J^{\mu}_{5}=\bar{q}_{i}\gamma_{5}\gamma^{\mu}q_{i}$ even under RI/MOM renormalization Pang:2024sdl.
+Pang et al.
+(2024), the ultraviolet (UV) divergence of $K^{\mu}$ is gauge independent.
+Therefore, we can perform a nonperturbative renormalization of $K^{\mu}$ in the $\mathrm{RI/MOM}$ scheme Martinelli et al.
+(1995) where the renormalization factor is typically computed in the Landau gauge.
+$K^{\mu}$ mixes only with the axial-vector current $J^{\mu}_{5}=\bar{q}\gamma_{5}\gamma^{\mu}q$ even under RI/MOM renormalization Pang et al.
+(2024).
 The renormalization constants can be determined by the following conditions
-$\begin{pmatrix}\langle ps|K^{\mu}|ps\rangle^{\mathrm{tree}}\\ \langle ps|J^{\mu}_{5}|ps\rangle^{\mathrm{tree}}\end{pmatrix}=\begin{pmatrix}Z_{11}^{\mathrm{RI}}&Z_{12}^{\mathrm{RI}}\\ Z_{21}^{\mathrm{RI}}&Z_{22}^{\mathrm{RI}}\end{pmatrix}\begin{pmatrix}\langle ps|K^{\mu}|ps\rangle^{\mathrm{lat.}}\\ \langle ps|J^{\mu}_{5}|ps\rangle^{\mathrm{lat.}}\end{pmatrix}$, (3)
-where $|ps\rangle$ denotes either gluon ($g$) or quark ($q$) external states with specific momentum and polarization.
-In the discussion below, we introduce the abbreviations $\mathcal{O}^{\mathrm{tree/lat.},{g/q}}\equiv\langle{g,q}|\mathcal{O}|{g,q}\rangle^{\mathrm{tree/lat.}}$ to simplify the notation.
+$\begin{pmatrix}K^{\mathrm{tree},g/q}\\ J_{5}^{\mathrm{tree},g/q}\end{pmatrix}=\begin{pmatrix}Z_{11}^{\mathrm{RI}}&Z_{12}^{\mathrm{RI}}\\ Z_{21}^{\mathrm{RI}}&Z_{22}^{\mathrm{RI}}\end{pmatrix}\begin{pmatrix}K^{\mathrm{lat.},g/q}\\ J_{5}^{\mathrm{lat.},g/q}\end{pmatrix}$, (3)
+where we introduce the abbreviations $\mathcal{O}^{\mathrm{tree/lat.},{g/q}}\equiv\langle{g/q}|\mathcal{O}|{g/q}\rangle^{\mathrm{tree/lat.}}$ to simplify the notation, where $|g\rangle$ or $|q\rangle$ denotes either gluon or quark external states with specific momentum and polarization.
 Note that $K^{\mathrm{tree},q}=J_{5}^{\mathrm{tree},g}=0$.
-Moreover, lattice data indicate that $J_{5}^{\mathrm{lat.},g}\approx 0$ within statistical uncertainties.
-Thus, the renormalization constants can be determined as:
+Moreover, since only disconnected quark loops are involved, we have $J_{5}^{\mathrm{lat.},g}\approx 0$.
+Thus, the RI/MOM renormalization constants can be determined as:
 $Z_{11}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\frac{K^{\mathrm{tree,g}}}{K^{\mathrm{lat.,g}}},\quad\quad Z_{12}^{\mathrm{RI}}(\mu_{\rm RI}^{2})=-\frac{K^{\mathrm{lat.,q}}}{J_{5}^{\mathrm{lat.,q}}}Z_{11}^{\mathrm{RI}},$
@@
 With the above approximation, $Z_{22}^{\mathrm{RI}}$ reduces to the axial-vector current renormalization constant $Z_{A}^{\mathrm{RI}}$ in Ref.
-Hu:2023jet.
+Hu and others (2024).
 The ratio $K^{\mathrm{lat.,q}}/J_{5}^{\mathrm{lat.,q}}$ appearing in the off-diagonal element $Z_{12}^{\mathrm{RI}}$ is of $\mathcal{O}(\alpha_{s})$.
@@
 $\Delta G^{\mathrm{B.}}$ and $\Delta\Sigma^{\mathrm{B.}}$ denote the bare gluon and quark helicities extracted from the bare matrix elements of $K^{\mu}$ and $J_{5}^{\mu}$, respectively.
-The RI/MOM renormalized results can then be converted to the $\overline{\mathrm{MS}}$ scheme via Pang:2024sdl
+The RI/MOM renormalized results can then be converted to the $\overline{\mathrm{MS}}$ scheme via Pang et al. (2024)
 $\Delta G^{\overline{\mathrm{MS}}}$ $=$ $R_{11}\Delta G^{\mathrm{RI}}+R_{12}\Delta\Sigma^{\mathrm{RI}},$
@@
 In the present work, we have included the conversion factors up to three-loops.
-Their detailed expressions are given in the Supplemental Material supplemental.
-Lattice Calculation: Our calculation is carried out on the $n_{f}=$2+1 flavor full QCD ensembles generated by CLQCD using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action Hu:2023jet.
+Their detailed expressions are given in the Supplemental Material A 1.
+Lattice Calculation: Our calculation is carried out on the $n_{f}=$2+1 flavor full QCD ensembles generated by CLQCD using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action Hu and others (2024); Du and others (2025).
 The ensembles used for the simulation are described in Table 1.
-The values of $m_{\pi}$ of all the ensembles are chosen to ensure $m_{\pi}L\geq 3.7$, thereby efficiently suppressing the finite-volume effect.
-For more details, we refer the readers to the Supplemental Material supplemental.
+The values of $m_{\pi}$ of all the ensembles are chosen to ensure $m_{\pi}L\geq 3.7$, thereby efficiently suppressing finite-volume effects.
+For more details, we refer the readers to the Supplemental Material B 1.
 Symbol $\hat{\beta}$ $a$ (fm) $\tilde{L}^{3}\times\tilde{T}$ $m_{\pi}$ (MeV) $N_{\mathrm{cfg}}$
 C24P29 6.200 0.10524(05)(62) $24^{3}\times 72$ 292.3(1.0) 780
-C48P23 6.200 0.10524(05)(62) $48^{3}\times 96$ 224.1(1.2) 400
+C48P23 $48^{3}\times 96$ 224.1(1.2) 400
 E32P29 6.308 0.08973(20)(53) $32^{3}\times 64$ 287.3(2.5) 890
 F32P30 6.410 0.07753(03)(45) $32^{3}\times 96$ 300.4(1.2) 800
-Table 1: Gauge coupling $\hat{\beta}=10/(g_{0}^{2}u_{0}^{4})$, lattice spacing $a$, dimensionless lattice size $\tilde{L}^{3}\times\tilde{T}$, corresponding pion mass $m_{\pi}$, and number of configurations $N_{\mathrm{cfg}}$ used for the simulation.
+Table 1: Lattice setup used for the simulation, with gauge coupling $\hat{\beta}=10/(g_{0}^{2}u_{0}^{4})$, lattice spacing $a$, dimensionless lattice size $\tilde{L}^{3}\times\tilde{T}$, corresponding pion mass $m_{\pi}$, and number of configurations $N_{\mathrm{cfg}}$.
 In applying Eq.
@@
 Furthermore, 5-HYP smearing is applied to the topological current throughout this calculation.
-For proton external states at low momenta ($p\leq 0.5\ \mathrm{GeV}$), we implement the distillation smearing technique HadronSpectrum:2009krc, while for higher momentum states we employ an approach combining distillation with momentum smearing Bali:2016lva.
-We insert the topological current operator into the two-point function of a proton polarized along the $i$th-direction (denoted as $\mathrm{Proj.i}$) to obtain the disconnected three-point function
-$C_{3}(t_{f},t_{i},t_{s})\equiv\langle\mathrm{PS}_{\mathrm{Proj.i}}(t_{f})|K^{t/i}_{\mathrm{C.G.}}(t_{i})|\mathrm{PS}_{\mathrm{Proj.i}}(t_{s})\rangle,$ (7)
-where $t_{f}$, $t_{i}$, and $t_{s}$ correspond to sink, current insertion, and source time-slices, respectively.
+For proton external states at low momenta ($p\leq 0.5\ \mathrm{GeV}$), we implement the distillation smearing technique Peardon et al.
+(2009), while for higher momentum states we employ an approach combining distillation with momentum smearing Bali et al.
+(2016); Egerer et al.
+(2021).
+We insert the topological current operator into the two-point function of a proton polarized along the $i$th-direction to obtain the disconnected three-point function $C_{3}(t_{f},t_{i},t_{s})$, where $t_{f}$, $t_{i}$, and $t_{s}$ correspond to sink, current insertion, and source time-slices, respectively.
 The bare matrix elements (BMEs) of the topological current are extracted through the following ratio:
 $R_{K^{\mu}}(t_{f},t_{i},t_{s})=\frac{C_{3}(t_{f},t_{i},t_{s})}{C_{2}(t_{f},t_{s})},$ (7)
-with $C_{2}(t_{f},t_{s})\equiv\langle\mathrm{PS}_{\mathrm{UnProj.}}(t_{f})|\mathrm{PS}_{\mathrm{UnProj.}}(t_{s})\rangle$ representing the unpolarized proton two-point correlation function.
-To optimize signal quality, we perform a summation over all source time-slices: $R_{K^{\mu}}(t_{f},t_{i})=\sum_{t_{s}}R_{K^{\mu}}(t_{f},t_{i},t_{s})$, which is then fitted to the form
-$R_{K^{\mu}}(t_{f},t_{i})=\langle K^{\mu}\rangle+c_{1}e^{-\Delta E(t_{f}-t_{i})}+c_{2}e^{-\Delta Et_{i}},$ (9)
-where $\langle K^{\mu}\rangle$ serves as an abbreviation for $\langle\mathrm{PS}_{\mathrm{Proj.i}}|K^{\mu}|\mathrm{PS}_{\mathrm{Proj.i}}\rangle/(2E)$, and $c_{1},c_{2},\Delta E$ are fitting parameters.
+with $C_{2}(t_{f},t_{s})$ representing unpolarized proton two-point correlation functions.
+We sum over all source time-slices: $R_{K^{\mu}}(t_{f},t_{i})=\sum_{t_{s}}R_{K^{\mu}}(t_{f},t_{i},t_{s})$, and fit the result to the form
+$R_{K^{\mu}}(t_{f},t_{i})=\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}+c_{1}e^{-\Delta E(t_{f}-t_{i})}+c_{2}e^{-\Delta Et_{i}},$ (8)
+where $\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}$ is defined as nuclear BME divided by the field operator’s normalization factor $2E$, and $c_{1},c_{2},\Delta E$ are fitting parameters.
 For illustrative purposes, we show in FIG.
-1 the fitted ground state matrix element for both $K^{i}$ and $K^{t}$ at $p=0.98$ GeV.
-The fitting results for other momenta can be found in the Supplemental Material supplemental.
+1 the fitted ground state BME for both $K_{i}$ and $K_{t}$ at $p=0.98$ GeV.
+The fitting results for other momenta can be found in the Supplemental Material B 1.
 Refer to caption Refer to caption
-Figure 1: The ratio $R_{K^{\mu}}(t_{f},t_{i})$ at $t_{f}\in[0.63,1.16]\ \mathrm{fm}$ and the fitted BME $\langle K^{\mu}\rangle$ (the gray error band) for proton external states with p=0.98GeVp=0.98\ $\mathrm${GeV} under ensemble C24P29, utilizing the fitting procedure described in (9).
+Figure 1: The ratio $R_{K^{\mu}}(t_{f},t_{i})$ at $t_{f}\in[0.63,1.16]\ \mathrm{fm}$ and the fitted BME $\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}$ (the gray error band) for proton external states with p=0.98GeVp=0.98\ $\mathrm${GeV} and ensemble C24P29, utilizing the fitting procedure described in Eq.
+(8).
 To obtain the renormalization constants, we evaluate the bare Green’s function for the topological current
@@
 The bare gluon propagator is given by
-$S_{g}^{\mu\nu}(p)=\mathrm{CV}\langle\mathrm{Tr}[A^{\mu}(p)A^{\nu}(-p)]\rangle,$ (11)
-where $\mathrm{CV}=2/[(N^{2}_{c}-1)\tilde{V}$], with $N_{c}=3$ representing the number of colors and $\tilde{V}=\tilde{L}^{3}\times\tilde{T}$ denoting the dimensionless lattice volume.
-Due to rotational symmetry breaking on the lattice, we can define the gluon field renormalization constants through two distinct prescriptions:
-$Z_{g,\mathrm{diag}}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\frac{\sum_{p_{\rho}=0}1}{\hat{p}^{2}\sum_{p_{\rho}=0}S_{g}^{\rho\rho}(p^{2})},$
-$Z_{g,\mathrm{off}}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\frac{\sum_{p_{\rho}\neq 0,p_{\tau}\neq 0,\rho\neq\tau}1}{\hat{p}^{4}\sum_{p_{\rho}\neq 0,p_{\tau}\neq 0,\rho\neq\tau}\frac{S_{g}^{\rho\tau}(p^{2})}{\hat{p}_{\rho}\hat{p}_{\tau}}},$ (12)
+$S_{g}^{\mu\nu}(p)=\mathrm{C_{V}}\langle\mathrm{Tr}[A^{\mu}(p)A^{\nu}(-p)]\rangle,$ (10)
+where $\mathrm{C_{V}}=2/[(N^{2}_{c}-1)\tilde{V}$], with $N_{c}=3$ representing the number of colors and $\tilde{V}=\tilde{L}^{3}\times\tilde{T}$ denoting the dimensionless lattice volume.
+In addition, we can define the gluon field renormalization constant
+$Z_{g,\mathrm{diag}}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\frac{\sum_{p_{\rho}=0}1}{\hat{p}^{2}\sum_{p_{\rho}=0}S_{g}^{\rho\rho}(p^{2})},$ (11)
 where $\hat{p}_{\mu}=\frac{2}{a}\sin{\frac{p_{\mu}a}{2}}$ represents the lattice momentum, and $\mu_{\rm RI}^{2}=p^{2}$ denotes the RI/MOM renormalization scale.
-$Z_{11}^{\mathrm{RI}}$ in Eq.(Total Gluon Helicity Contribution to Proton Spin from Lattice QCD) then takes the following form
-$Z_{11}^{\mathrm{RI}}(\mu_{\rm RI}^{2})=\frac{Z_{g,\mathrm{diag/off}}(\mu_{\rm RI}^{2})}{\mathrm{Tr}[S_{g}^{-1}(p_{1})G_{K}(p_{1},p_{2})S_{g}^{-1}(p_{2})(K^{\mathrm{tree,g}})^{-1}]},$ (13)
+$Z_{11}^{\mathrm{RI}}$ in Eq.(Total Gluon Helicity Contribution to the Proton Spin from Lattice QCD) then takes the following form
+$Z_{11}^{\mathrm{RI}}(\mu_{\rm RI}^{2})=\frac{Z_{g,\mathrm{diag}}(\mu_{\rm RI}^{2})}{\mathrm{Tr}[S_{g}^{-1}(p_{1})G_{K}(p_{1},p_{2})S_{g}^{-1}(p_{2})(K^{\mathrm{tree,g}})^{-1}]},$ (12)
 where the external momenta are chosen to satisfy $p_{1}=p_{2}=p$.
 The Green’s function of $K^{\mathrm{tree,g}}$ reads $i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}$ in our normalization convention.
-In addition, we hide the indicator of the bare Green’s function $G_{K}$ and bare gluon propagator $S_{g}$ in (13).
-The inverse of $Z_{11}^{\rm RI}$ then reads,
-$Z_{11}^{-1,\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\sum_{\mu,\rho,\nu,\sigma}\frac{2\mathrm{CV}\hat{p}^{2}G_{K}^{\mu\rho\nu}(p,p)}{i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}(S_{g}^{\mu\mu}+S_{g}^{\nu\nu})(p^{2})}|_{\mu_{\rm RI}^{2}=p^{2}}$ (14)
-$=$ $\sum_{\mu,\rho,\nu,\sigma}\frac{\mathrm{CV}\hat{p}^{4}G_{K}^{\mu\rho\nu}(p,p)}{i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}\hat{p}_{\mu}\hat{p}_{\nu}S_{g}^{\mu\nu}(p^{2})}|_{\mu_{\rm RI}^{2}=p^{2}}.$
+We hide the $\mu\rho\nu$ indicator of the bare Green’s function $G_{K}$ and bare gluon propagator $S_{g}$ in Eq.
+(12).
+$Z_{11}^{\rm RI}$ then reads,
+$Z_{11}^{\mathrm{RI}}(\mu_{\rm RI}^{2})=\sum_{\mu,\rho,\nu,\sigma}\frac{i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}(S_{g}^{\mu\mu}+S_{g}^{\nu\nu})(p^{2})}{2\mathrm{C_{V}}\hat{p}^{2}G_{K}^{\mu\rho\nu}(p,p)}|_{\mu_{\rm RI}^{2}=p^{2}}.$ (13)
 The summation over indices $\mu,\nu,\rho,\sigma$ in conjunction with the Levi-Civita antisymmetric tensor yields a total of 12 distinct configurations.
-To enhance the signal quality of the pure gluon Green’s function $G_{K}(p,p)$ in (14), we employ the cluster decomposition error reduction (CDER) technique Liu:2017man, which requires large volume ensembles.
+To enhance the signal quality of the pure gluon Green’s function $G_{K}(p,p)$ in Eq.
+(13), we abandoned original brute force calculation (which requires huge statistics) and instead adopted the Cluster Decomposition Error Reduction (CDER) technique Liu et al.
+(2018), which requires a large volume ensemble.
 Due to the limited volume of C24P29, we have chosen to compute the renormalization constants on a larger volume ensemble, C48P23, which shares the same lattice spacing.
 In contrast, for the other two ensembles E32P29 and F32P30, the CDER can be directly applied.
-For more details on the CDER implementation, we refer the readers to the Supplemental Material supplemental.
-The conversion factors from the RI/MOM scheme to the $\overline{\rm MS}$ scheme are evaluated by choosing $\mu/\mu_{\rm RI}=e^{-{5}/{6}}$ Brodsky:1982gc in the fixed-order expression up to three-loops, and then RG-evolved to $\mu^{2}=10\,{\rm GeV}^{2}$.
+For more details on the CDER implementation, we refer readers to the Supplemental Material C 1.
+The conversion factors from the RI/MOM scheme to the $\overline{\rm MS}$ scheme are evaluated by choosing $\mu/\mu_{\rm RI}=e^{-{5}/{6}}$ Brodsky et al.
+(1983) in the fixed-order expression up to three-loops, and then RG-evolve to $\mu^{2}=10\,{\rm GeV}^{2}$.
 Subsequently, the discretization effects of ${\cal O}(a^{2}p^{2})$ are removed following the strategy proposed in Ref.
-Yang:2018nqn.
+Yang et al.
+(2018b).
 Note that here the extrapolation is conducted by varying $p^{2}$ (and thus $a^{2}p^{2}$) and determining the value at $a^{2}p^{2}=0$.
-To simultaneously ensure the reliability of RI/MOM renormalization and an acceptable signal-to-noise ratio after applying the CDER, we choose the data in the momentum interval $\mu_{\mathrm{RI}}\in[3.0,5.0]\ \mathrm{GeV}$ for the extrapolation, which correspond to $a^{2}p^{2}\in[2.6,7.1]$ under ensemble C48P23.
+To simultaneously ensure the reliability of RI/MOM renormalization and an acceptable signal-to-noise ratio after applying the CDER, we choose the data in the momentum interval $\mu_{\mathrm{RI}}\in[3.0,5.0]\ \mathrm{GeV}$ for the extrapolation, which correspond to $a^{2}p^{2}\in[2.6,7.1]$ for ensemble C48P23 (the CDER cutoff parameters for $K^{\mu}$ and $A^{\mu}$ are 1.26fm1.26\ $\mathrm${fm} and 0.74fm0.74\ $\mathrm${fm}, respectively).
 The extrapolated result for the inverse of $\bar{Z}_{11}^{\mathrm{lat}}\equiv R_{11}Z_{11}^{\mathrm{RI}}$ is shown in Fig.
@@
 Similar analyses have also been performed for the other two ensembles with different lattice spacings.
-Regarding the mixing with quark spin $\Delta\Sigma$, it turns out to be an order of magnitude less than $\bar{Z}_{11}^{\mathrm{lat}}$, especially at large $a^{2}p^{2}$ (See the Supplemental Material supplemental for more details).
+The mixing with quark spin $\Delta\Sigma$, turns out to be an order of magnitude less than $\bar{Z}_{11}^{\mathrm{lat}}$, especially at large $a^{2}p^{2}$ (See the Supplemental Material A 1 for more details).
 These properties offer a significant simplification in our numerical analysis, leading us to
 $\Delta G^{\overline{\mathrm{MS}}}\approx\bar{Z}_{11}^{\mathrm{lat}}\Delta G^{\mathrm{B.}},\quad\Delta\Sigma^{\overline{\mathrm{MS}}}\approx\bar{Z}_{A}^{\mathrm{lat}}\Delta\Sigma^{\mathrm{B.}}.$ (14)
-Figure 2: The removal of ${\cal O}(a^{2}p^{2})$ discretization effect in $\bar{Z}_{11}^{\mathrm{lat}}\equiv R_{11}Z_{11}^{\mathrm{RI}}$ for the ensemble C48P23.
-The red, green, and blue data points and their error bands represent the CDER results and their $a^{2}p^{2}\rightarrow 0$ extrapolated results using momentum modes $(0,0,p,p)$, $(0,p,p,p)$, and $(p,p,p,p)$ in (14), where $p=\big(2\pi/(aL)\big)p_{\mathrm{lat}}$ with $p_{\mathrm{lat}}$ represents momentum in the lattice unit.
+Figure 2: The removal of ${\cal O}(a^{2}p^{2})$ discretization effect in $\bar{Z}_{11}^{\mathrm{lat}}$ for the ensemble C48P23.
+The red, green, and blue data points and their error bands represent the CDER results and their $a^{2}p^{2}\rightarrow 0$ extrapolated results using momentum modes $(0,0,p,p)$, $(0,p,p,p)$, and $(p,p,p,p)$ in Eq.
+(13), where $p=\big(2\pi/(aL)\big)p_{\mathrm{lat}}$ with $p_{\mathrm{lat}}$ represents momentum in the lattice unit.
 Numerical Results: After obtaining the renormalized matrix elements of $K^{\mu}$ and converting them to the $\overline{\rm MS}$ scheme, we are able to extract the final result by carrying out a continuum and infinite momentum extrapolation.
 Fig.
-3 displays $\langle K^{\mu}\rangle$ as a function of lattice spacings and proton momenta.
+3 displays $\langle K^{\mu}\rangle_{N}\equiv\bar{Z}_{11}^{\mathrm{lat.}}\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}$ as a function of lattice spacings and proton momenta.
 As can be seen from the plot, the results for $\mu=t,i$ exhibit a clear trend consistent with Lorentz covariance within uncertainties.
 Motivated by this, we adopt the following form for the continuum and infinite momentum extrapolation:
-$\langle K^{\mu}\rangle=\frac{S^{\mu}}{E}\big(\Delta G+\frac{c_{\mathrm{h.t.}}}{E^{2}}\big)+c_{a}a^{2},$ (16)
+$\langle K^{\mu}\rangle_{N}=\frac{S^{\mu}}{E}\big(\Delta G+\frac{c_{\mathrm{h.t.}}}{E^{2}}\big)+c_{a}a^{2},$ (15)
 where $\Delta G$ is the result to be extracted, $c_{\mathrm{h.t.}}$ accounts for higher-twist contributions, and $c_{a}$ captures discretization effects.
 The pion mass dependence has been neglected in this analysis, primarily due to the significant computational cost of gluon matrix elements at the physical point and the very mild pion mass dependence observed in Ref.
-Yang:2016plb.
-Since the proton momenta are not significantly larger than its mass, we incorporate the higher-twist contribution using an empirical parametrization form $1/E^{2}$ in the equation above, where $E=\sqrt{M^{2}+p^{2}}$ with the proton mass $M\approx 1.02\ \mathrm{GeV}$ according to the dispersion relation.
-The primary sources of systematic uncertainty in our analysis are identified as follows: 1) Including or not the zero-momentum data points in the extrapolation; 2) Discretization errors associated with the finite lattice spacing; 3) The choice of the UV window in the RI/MOM scale or $a^{2}p^{2}$ (the interval $\mu_{\mathrm{RI}}\in[3.0,5.0]\ \mathrm{GeV}$ has been extended to [3.0,6.0]GeV[3.0,6.0]\ $\mathrm${GeV} to estimate the corresponding uncertainty); and 4) Systematic errors associated with the perturbative inputs (estimated by varying from three-loop to two-loop accuracy).
+Yang et al.
+(2017).
+Since the proton momenta are not significantly larger than its mass, we incorporate the higher-twist contribution using an empirical parametrization form $1/E^{2}$ in the equation above, where $E=\sqrt{M^{2}+p^{2}}$ with the proton mass $M\approx 1\ \mathrm{GeV}$ according to the dispersion relation.
+The primary sources of systematic uncertainty in our analysis are identified as follows: 1) Including or not the zero-momentum data points in the extrapolation; 2) Discretization errors associated with the finite lattice spacing (the central value difference between the zero momentum $\langle K_{i}\rangle_{N}$ for the finest lattice spacing and that extrapolated to the continuum limit); 3) The choice of the UV window in the RI/MOM scale or $a^{2}p^{2}$ (the interval $\mu_{\mathrm{RI}}\in[3.0,5.0]\ \mathrm{GeV}$ has been extended to [3.0,6.0]GeV[3.0,6.0]\ $\mathrm${GeV} to estimate the corresponding uncertainty); and 4) Systematic errors associated with the perturbative input (estimated by varying from three-loop to two-loop accuracy).
 As illustrated in FIG.
 3, the joint extrapolation yields a gluon helicity contribution $\Delta G^{\overline{\mathrm{MS}},\ 10\ \mathrm{GeV}^{2}}=0.231(17)^{\mathrm{sta.}}(44)^{\mathrm{sym.}}$, which constitutes $46(9)\%$ of the proton’s total spin.
-Figure 3: The continuum and infinite momentum extrapolation of the total gluon helicity results using lattice spacings ranging from 0.08fm0.08\ $\mathrm${fm} to 0.11fm0.11\ $\mathrm${fm} at the $\overline{\mathrm{MS}}$ scale $\mu^{2}=10\ \mathrm{GeV}^{2}$.
+Figure 3: The continuum and infinite momentum extrapolation of the total gluon helicity is performed using lattice spacings from 0.08fm0.08\ $\mathrm${fm} to 0.11fm0.11\ $\mathrm${fm} at the $\overline{\mathrm{MS}}$ scale $\mu^{2}=10\ \mathrm{GeV}^{2}$.
 The uncertainties are shown as separate bands: statistical uncertainty (dark gray) and all uncertainties that arise during our analysis (light gray).
@@
 The result is then extrapolated to the continuum and infinite momentum limits, yielding $\Delta G=0.231(17)^{\mathrm{sta.}}(44)^{\mathrm{sym.}}$ at $\mu^{2}=10\ \mathrm{GeV}^{2}$.
-This corresponds to approximately $46(7)\%$ of the proton spin and clearly rules out the possibility of a negative gluon helicity contribution Zhou:2022wzm; Whitehill:2022mpq; Hunt-Smith:2024khs.
+This corresponds to approximately $46(9)\%$ of the proton spin and clearly rules out the possibility of a negative gluon helicity contribution Zhou et al.
+(2022); Whitehill et al.
+(2023); Hunt-Smith et al.
+(2024).
 The finding offers important insight for future theoretical and experimental efforts aimed at resolving the proton spin puzzle.
@@
 We thank Fangcheng He, Zhi-Cheng Hu, Xiangyu Jiang, Li-Jun Zhou, and other CLQCD and LPC members for valuable comments and suggestions.
-We thank the CLQCD collaborations for providing us their gauge configurations with dynamical fermions Hu:2023jet, which are generated on HPC Cluster of ITP-CAS, IHEP-CAS and CSNS-CAS, the Southern Nuclear Science Computing Center(SNSC) and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University.
-The calculations were performed using the PyQUDA Jiang:2024lto and Chroma Edwards:2004sx software suite with QUDA Clark:2009wm; Babich:2011np; Clark:2016rdz through HIP programming model Bi:2020wpt.
+We thank the CLQCD collaborations for providing us their gauge configurations with dynamical fermions Hu and others (2024), which are generated on HPC Cluster of ITP-CAS, IHEP-CAS and CSNS-CAS, the Southern Nuclear Science Computing Center(SNSC) and the Siyuan-1 cluster supported by the Center for High Performance Computing at Shanghai Jiao Tong University.
+The calculations were performed using the PyQUDA Jiang et al.
+(2024) and Chroma Edwards and Joo (2005) software suite with QUDA Clark et al.
+(2010); Babich et al.
+(2011); Clark et al.
+(2016) through HIP programming model Bi et al.
+(2020).
 The numerical calculation were carried out on the ORISE Supercomputer, the Southern Nuclear Science Computing Center (SNSC).
@@
 tsqn202312052 and 2024HWYQ-005.
-Supplemental Material
-.1 Detailed information of CLQCD ensembles
-The results in this work, are based on the $n_{f}=2+1$ flavor ensembles from the CLQCD collaboration using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action Hu:2023jet.
-The TITLS gauge action is defined in the following,
-$S_{g}=\frac{1}{N_{c}}\mathrm{Re}\sum_{x,\mu<\nu}\mathrm{Tr}\left[1-\hat{\beta}\left(\mathcal{P}^{U}_{\mu,\nu}(x)+\frac{c_{1}\mathcal{R}^{U}_{\mu,\nu}(x)}{1-8c_{1}^{0}}\right)\right],$ (17)
-and
-$\mathcal{P}^{U}_{\mu,\nu}(x)$ $=U_{\mu}(x)U_{\nu}(x+a\hat{\mu})U^{\dagger}_{\mu}(x+a\hat{\nu})U^{\dagger}_{\nu}(x),$
-$\mathcal{R}^{U}_{\mu,\nu}(x)$ $=U_{\mu}(x)U_{\mu}(x+a\hat{\mu})U_{\nu}(x+2a\hat{\mu})U^{\dagger}_{\mu}(x+a\hat{\mu}+a\hat{\nu})U^{\dagger}_{\mu}(x+a\hat{\nu})U^{\dagger}_{\nu}(x),$
-$U_{\mu}(x)$ $=P\left[\mathrm{exp}\left({\rm i}g_{0}\int_{x}^{x+\hat{\mu}a}\mathrm{d}yA_{\mu}(y)\right)\right],$
-$\hat{\beta}=(1-8c_{1}^{0})\frac{6}{g_{0}^{2}u_{0}^{4}}\equiv 10/(g_{0}^{2}u_{0}^{4})$ with $c_{1}^{0}=-\frac{1}{12}$, $c_{1}=\frac{c_{1}^{0}}{u_{0}^{2}}$, $u_{0}=\langle\frac{\mathrm{Re}\mathrm{Tr}\sum_{x,\mu<\nu}\mathcal{P}^{U}_{\mu\nu}(x)}{6N_{c}\tilde{V}}\rangle^{1/4}$ is the tadpole improvement factor.
-The TITLC fermion action uses 1-step stout smeared link $V$ with smearing parameter $\rho=0.125$,
-$S_{q}(m)=\sum_{x,\mu=1,...,4,\eta=\pm}\bar{\psi}(x)\sum\frac{1+\eta\gamma_{\mu}}{2}V_{\eta\mu}(x)\psi(x{+\eta\hat{\mu}a})$
-$\quad+\sum_{x}\psi(x)\left[-(4+ma)\delta_{y,x}+c_{\rm sw}\sigma^{\mu\nu}g_{0}F_{\mu\nu}\right]\psi(x),$ (18)
-where $c_{\rm sw}=\frac{1}{v^{3}_{0}}$ with $v_{0}=\langle\frac{\mathrm{Re}\mathrm{Tr}\sum_{x,\mu<\nu}\mathcal{P}^{V}_{\mu\nu}(x)}{6N_{c}\tilde{V}}\rangle^{1/4}$, and $F_{\mu\nu}$ has the same form as (21).
-The parameters used for the simulation are described in Table 1.
-.2 Bare nucleon matrix elements with topological current in the moving frame
-In lattice QCD computations of the topological current,
-$K^{\mu}(x)=\epsilon^{\mu\nu\rho\sigma}\text{Tr}[A_{\nu}F_{\rho\sigma}-\frac{2}{3}ig_{0}A_{\nu}A_{\rho}A_{\sigma}](x),$ (19)
-the fundamental components requiring evaluation are the gauge potential $A_{\mu}$ and the field strength tensor $F_{\mu\nu}$, defined through the following discretized formulations:
-$A_{\mu}(x+\hat{\mu}/2)=\left[\frac{U_{\mu}(x)-U_{\mu}^{\dagger}(x)}{2ig_{0}a}\right]_{\mathrm{traceless}.},$ (20)
-$F_{\mu\nu}$ $=$ $\frac{i}{8g_{0}a^{2}}(\mathcal{P}_{\mu,\nu}-\mathcal{P}_{\nu,\mu}+\mathcal{P}_{\nu,-\mu}-\mathcal{P}_{-\mu,\nu}$ (21)
-$+\mathcal{P}_{-\mu,-\nu}-\mathcal{P}_{-\nu,-\mu}+\mathcal{P}_{-\nu,\mu}-\mathcal{P}_{\mu,-\nu}),$
-where $U_{\mu}(x)$ denotes the gauge link variable and $\mathcal{P}_{\mu,\nu}=U_{\mu}(x)U_{\nu}(x+a\hat{\mu})U^{\dagger}_{\mu}(x+a\hat{\nu})U^{\dagger}_{\nu}(x)$ represents the plaquette operator.
-An alternative formulation employs the full-lattice definition of the gauge potential, $\tilde{A}_{\mu}(x)=(A_{\mu}(x+\hat{\mu}/2)+A_{\mu}(x-\hat{\mu}/2))/2$.
-The distinction between half-lattice and full-lattice definitions manifests primarily as different discretization errors.
-Consequently, we adopt the half-lattice definition of the gauge potential for all subsequent calculations.
-Based on the definition of local topological current in (19), one can give the expression of $K^{\mu}$ in different components, namely
-$K^{i}$ $\equiv$ $K^{i}_{E\times A}+K^{i}_{A_{t}B}+K^{i}_{3A}=2\text{Tr}[A\times E]+\epsilon^{tjk}\text{Tr}[A_{t}B_{k}]-\frac{2}{3}ig_{0}\epsilon^{jkt}\mathrm{Tr}[A_{j}A_{k}A_{t}],$
-$K^{t}$ $\equiv$ $K^{t}_{B\cdot A}+K^{t}_{3A}=\text{Tr}[A\cdot B]-\frac{2}{3}ig_{0}\epsilon^{ijk}\mathrm{Tr}[A_{i}A_{j}A_{k}],$ (22)
-where $E_{i}=F_{ti}$ and $B_{i}=F_{jk}$.
-The temporal component $K^{t}$, being predominantly determined by the magnetic field strength term $K^{t}_{B\cdot A}$, exhibits matrix elements that are consistent with zero within statistical uncertainties at zero momentum, yet demonstrate a gradual increase followed by saturation to a constant value with rising momentum.
-Conversely, the spatial components $K^{i}$, which are primarily governed by the electric field strength term $K^{i}_{E\times A}$, maintain approximately constant matrix elements within error margins across different momentum values.
-The physical interpretation for this behavior aligns with (16), stemming from the treatment of protons as point-like entities: the charge density represents an intrinsic property independent of momentum, while the current density evolves from zero at rest toward a limiting value with increasing momentum.
-These characteristic momentum dependencies constitute a direct manifestation of Lorentz covariance within a specific reference frame.
-To establish the presence of Lorentz covariance, we compute the matrix elements of the topological current operator for proton external states with varying momenta.
-This investigation requires two complementary computational stages: first, obtaining a stable plateau in the effective mass extraction from proton two-point correlation functions; second, calculating the disconnected three-point correlation functions with topological current insertions.
-Within the lattice QCD framework, appropriate smearing techniques must be employed to achieve adequate signal-to-noise ratios for both proton mass determinations and $K^{\mu}$ matrix element extractions.
-FIG.
-4 presents the proton effective mass behavior under ensemble C24P29.
-Analysis reveals that while the distillation smearing technique alone provides well-optimized plateau regions for the effective mass at low momenta, this stability deteriorates significantly at higher momenta (e.g., p=1.96GeVp=1.96\ $\mathrm${GeV}).
-In comparison, implementation of distillation + momentum smearing scheme yields moderately improved plateau stability at large momenta relative to the distillation-only approach.
-Refer to caption Refer to caption
-Figure 4: Proton effective mass with different momentum under different smearing schemes under ensemble C24P29.
-Left figure represents the distillation smearing scheme alone, and the right figure represents the distillation + momentum smearing.
-The red, green and blue points represent p=0,0.98,1.96GeVp=0,0.98,1.96\ $\mathrm${GeV}, respectively.
-Employing the identical smearing methodology as proton effective mass, we extract the BMEs of $K^{t/i}$ under ensemble C24P29.
-Ratio data presented in FIG.
-5 serve to validate the theoretical discussion surrounding (.2).
-In summary, the employed smearing technique demonstrates robust performance throughout the computational methodology.
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Figure 5: Ratio $R_{K^{\mu}}(t_{f},t_{i})$ at $t_{f}\in[0.63,1.16]\ \mathrm{fm}$ and the BMEs of $K^{\mu}/(2E)$ (the gray error band) for proton external states with varying momenta under ensemble C24P29, utilizing the fitting procedure described in (9).
-The results are organized such that the first column corresponds to spatial components $K^{i}$, while the second column represents temporal component $K^{t}$.
-Rows one through three display results for proton momenta of p=0,0.98,1.96GeVp=0,0.98,1.96\ $\mathrm${GeV}, respectively.
-.3 Non-perturbative renormalization with cluster decomposition error reduction (CEDR)
-The computation of the renormalization constant $Z_{11}^{\mathrm{RI}}$ within the $\mathrm{RI/MOM}$ scheme, as defined in (14), necessitates the evaluation of the three-point correlation function specified in (10).
-This calculation presents significant numerical challenges, even with precise determinations of the gluon propagator (11) at better than $1\%$ accuracy level.
-The extent of this difficulty is demonstrated in FIG.
-2, where the light background bands represent direct computations of $Z_{11}^{\mathrm{RI}}$ based on approximately 400 configurations of the $n_{f}=2+1$ TITLC + TITLS gauge ensemble C48P23 with lattice spacing $a\approx 0.1$ fm Hu:2023jet.
-These direct calculations exhibit substantial statistical uncertainties, rendering $Z_{11}^{\mathrm{RI}}$ effectively unresolved across all scales.
-Applying the cluster decomposition error reduction (CDER) technique enables significant error suppression in $Z_{11}^{\mathrm{RI}}$ calculations Liu:2017man; Yang:2018bft.
-The fundamental principle of cluster decomposition asserts that correlators exhibit exponential decay with increasing separation between operator insertions, implying that integration beyond the correlation length primarily contributes to statistical noise rather than meaningful signal.
-The CDER methodology addresses this by truncating the volume integral beyond a characteristic distance, thereby achieving an enhancement of the signal-to-noise ratio by a factor of $\sqrt{V}$, as demonstrated in Liu:2017man.
-When implementing CDER for the computation in (10), we introduce two distinct cutoff parameters: $r_{1}$ between the gluon operator and one gauge field, and $r_{2}$ between the gauge fields within the gluon propagator.
-This implementation yields the modified correlator with appropriate spatial cutoffs.
-The CDER technique achieves computational efficiency of $\mathcal{O}(V\log V)$ through the repeated application of fast Fourier transforms (FFTs).
-For the three-point function specified in (10), we implement the following computational strategy:
-1) Construct the spatially-restricted topological current operator: We compute $K^{\rho}_{r^{\prime}}(x)=\int_{|r^{\prime}|<r_{2}}d^{4}r^{\prime}K^{\rho}(x+r^{\prime})$ by first performing Fourier transforms (FT) of both $K^{\rho}(x)$ and the Heaviside function $f(x)=\theta(r_{2}-|x|)$, multiplying the transformed functions in momentum space, and subsequently applying anti-FT.
-2) Form the intermediate operator product: We calculate $B^{\rho}_{\mu,r^{\prime}}(x)=A_{\mu}(x)K^{\rho}_{r^{\prime}}(x)$ through direct multiplication in coordinate space.
-3) Apply cluster decomposition to the correlation function: For the expression $\int d^{4}xd^{4}ye^{ip\cdot(x-y)}B^{\rho}_{\mu,r^{\prime}}(x)A_{\nu}(y)$, we perform FT of both $A_{\nu}$ and $B^{\rho}_{\mu,r^{\prime}}$, apply anti-FT to the product $A_{\nu}(p)B^{\rho}_{\mu,r^{\prime}}(-p)$, implement an additional Heaviside function $g(x)=\theta(r_{1}-|x|)$ in coordinate space, and FT the resulting product.
-4) Get a cutoff bare Green’s function in this way:
-$G_{K}^{\mu\rho\nu,\mathrm{CDER}}(p,p)=\langle\int_{|r|<r_{1}}d^{4}r\int_{|r^{\prime}|<r_{2}}d^{4}r^{\prime}\int d^{4}xe^{ip\cdot r^{\prime}}K^{\rho}(x+r^{\prime})\mathrm{Tr}[A_{\mu}(x)A_{\nu}(x+r)]\rangle.$ (23)
-Here both momentum arguments of bare Green’s function are set to $p$ in accordance with the standard implementation of the $\mathrm{RI/MOM}$ scheme, where the renormalization scale is defined by $\mu_{\rm RI}^{2}=p^{2}$.
-The selection of cutoff parameters $r_{1,2}$ in the CDER technique requires careful consideration, as inappropriate choices may induce periodic oscillations in the renormalization constants Liu:2017man.
-FIG.
-6 provides visual confirmation of this analysis.
-When the cutoff parameter governing the gluon field-operator correlation ($r_{1}$) is set too small (top left panel), distinct periodic oscillations become apparent in the results.
-These oscillations progressively stabilize as $r_{1}$ increases.
-A similar pattern is observed for variations in the current cutoff parameter ($r_{2}$), though changes in oscillatory behavior is less pronounced without corresponding adjustments to $r_{1}$.
-In FIG.
-2, we employ the optimal cutoffs identified in the bottom right panel of FIG.
-6, specifically $r_{1}=1.26\ \mathrm{fm}$ and $r_{2}=0.74\ \mathrm{fm}$, to perform discretization error subtraction procedure.
-Thus, we have successfully computed the purely gluonic Green’s function using the CDER technique, which is crucial for the non-perturbative renormalization of topological current.
-This method effectively suppresses the significant statistical noise generated when directly computing the three-point correlation function by truncating the volume integral beyond the characteristic distances $r_{1}$ and $r_{2}$, thus enabling the stable determination of the renormalization constant $Z_{11}^{\mathrm{RI}}$, which was previously only obtainable through the stacking statistics.
-Refer to caption Refer to caption
-Refer to caption Refer to caption
-Figure 6: The $\overline{\mathrm{MS}}$ renormalization constant $\bar{Z}_{11}^{\mathrm{lat}}(a^{2}p^{2},{\mu}^{2}=10\ \mathrm{GeV}^{2})$ under ensemble C48P23 as a function of $a^{2}p^{2}$ using momentum modes $(p_{\mathrm{lat}},p_{\mathrm{lat}},p_{\mathrm{lat}},p_{\mathrm{lat}})$, with different cutoffs on the gluon field-operator correlation ($r_{1}$) and current ($r_{2}$).
-.4 Three-loop approximation of $\Delta G$’s off-diagonal elements
-This subsection outlines the derivation of the reduced expressions presented in (15).
-Starting from (Total Gluon Helicity Contribution to Proton Spin from Lattice QCD) and (Total Gluon Helicity Contribution to Proton Spin from Lattice QCD), we obtain the exact formulations for gluon helicity in the $\overline{\mathrm{MS}}$ scheme:
-$\Delta G^{\overline{\mathrm{MS}}}(a^{2}p^{2},{\mu}^{2})$ $=$ $R_{11}(a^{2}p^{2},{\mu}^{2})Z_{11}^{\mathrm{RI}}(a^{2}p^{2})\big(\Delta G^{\mathrm{B.}}-\frac{K^{\mathrm{lat.,q}}}{J_{5}^{\mathrm{lat.,q}}}(a^{2}p^{2})\Delta\Sigma^{\mathrm{B.}}\big)$ (24)
-$+R_{12}(a^{2}p^{2},{\mu}^{2})Z_{A}^{\mathrm{RI}}(a^{2}p^{2})\Delta\Sigma^{\mathrm{B.}}.$
-In this formulation, we express the dependence on the squared renormalization scale $\mu_{\rm RI}^{2}$ of the $\mathrm{RI/MOM}$ scheme in terms of the dimensionless quantity $a^{2}p^{2}$, and the same below.
-We now implement a three-loop approximation for the ratio term $K^{\mathrm{lat,q}}/{J_{5}^{\mathrm{lat,q}}}$.
-Noting that $R_{12}$ can be determined from the topological current $K^{\mu}$ under quark external states while $R_{22}\approx 1$ is fixed by the axial vector current under quark external states, we employ the approximation:
-$\big(\bar{Z}_{11}^{\mathrm{lat}}R_{12}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2})$ $\approx$ $\big(\bar{Z}_{11}^{\mathrm{lat}}\frac{R_{12}}{R_{22}}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2})$ (25)
-$=$ $\big(\bar{Z}_{11}^{\mathrm{lat}}\frac{K^{3\mathrm{-loop,q}}}{J_{5}^{3\mathrm{-loop,q}}}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2})$
-$\approx$ $\big(\bar{Z}_{11}^{\mathrm{lat}}\frac{K^{\mathrm{lat,q}}}{J_{5}^{\mathrm{lat,q}}}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2}).$
-where we used a redefined expression $\bar{Z}_{11}^{\mathrm{lat}}\equiv R_{11}Z_{11}^{\mathrm{RI}}$.
-Furthermore, owing to the protection by chiral symmetry, the matching coefficient for the axial-vector current remains unity up to two loops, i.e., $\bar{Z}_{A}^{\mathrm{lat}}\equiv Z_{A}^{\overline{\mathrm{MS}}}\approx Z_{A}^{\mathrm{RI}}$.
-This convention will be consistently maintained throughout the subsequent discussion.
-This approximation enables the following simplified expression for gluon helicity at three-loop order:
-$\Delta G^{\overline{\mathrm{MS}}}$ $\overset{3\mathrm{-loop}}{\approx}$ $\bar{Z}_{11}^{\mathrm{lat}}\big(\Delta G^{\mathrm{B.}}-R_{12}\Delta\Sigma^{\mathrm{B}.}\big)+R_{12}\bar{Z}_{A}^{\mathrm{lat}}\Delta\Sigma^{\mathrm{B.}}$ (26)
-$=$ $\bar{Z}_{11}^{\mathrm{lat}}\Delta G^{\mathrm{B.}}+(\bar{Z}_{A}^{\mathrm{lat}}-\bar{Z}_{11}^{\mathrm{lat}})R_{12}\Delta\Sigma^{\mathrm{B.}}$
-$\approx$ $\bar{Z}_{11}^{\mathrm{lat}}\Delta G^{\mathrm{B.}},$
-where the final reduction is justified based on the hierarchy $\bar{Z}_{11}^{\mathrm{lat}}\gg(\bar{Z}_{A}^{\mathrm{lat}}-\bar{Z}_{11}^{\mathrm{lat}})R_{12}$, and by the fact that $\Delta G^{\mathrm{B.}}$ is comparable in magnitude to $\Delta\Sigma^{\mathrm{B.}}$.
-The derivation of expression $\Delta\Sigma^{\overline{\mathrm{MS}}}\approx\bar{Z}_{A}^{\mathrm{lat}}\Delta\Sigma^{\mathrm{B.}}$ parallels the approach used in the approximation above.
-.5 Perturbative results for matching coefficients $R_{ij}$ up to three-loops
-This section details the perturbative calculation of the finite renormalization (or matching) coefficients $R_{ij}$ introduced in eq.
-(Total Gluon Helicity Contribution to Proton Spin from Lattice QCD) for the singlet axial-vector current $J_{5}^{\mu}$ and topological current $K^{\mu}$ from the intermediate $\mathrm{RI/MOM}$ scheme used in Lattice calculation to the $\overline{\mathrm{MS}}$ scheme typically employed in perturbative calculations in Dimensional Regularization (DR).
-The explicit results for $R_{ij}$ up to three loops are documented for the sake of references.
+Appendix A Supplemental Material
+A.1 A. Perturbative results for matching coefficients $R_{ij}$ up to three-loops
+According to Ref.
+Pang et al.
+(2024), converting the singlet axial-vector current $J_{5}^{\mu}$ and the topological current $K^{\mu}$ from the lattice $\mathrm{RI/MOM}$ scheme to the continuum $\overline{\mathrm{MS}}$ scheme requires a finite renormalization (matching) coefficient $R_{ij}$.
+This section details the perturbative calculation of $R_{ij}$ in dimensional regularization.
+Extending the results of Ref.
+Pang et al.
+(2024), we provide explicit expressions for $R_{ij}$ up to three-loop order.
+Note that our result for the one-loop perturbative matching coefficient differs from that reported in Ref.
+Pang et al.
+(2024), with the differences implied in the following discussion.
+Furthermore, we demonstrate that the off-diagonal elements of the (matching) coefficient can be neglected, leading to a simplified matching relation.
+Following Ref. Pang et al. (2024), the following conditions apply within the RI/MOM renormalization scheme, namely
+$\begin{pmatrix}K^{\mathrm{tree},g/q}\\ J_{5}^{\mathrm{tree},g/q}\end{pmatrix}=\begin{pmatrix}Z_{11}^{\mathrm{RI}}&Z_{12}^{\mathrm{RI}}\\ Z_{21}^{\mathrm{RI}}&Z_{22}^{\mathrm{RI}}\end{pmatrix}\begin{pmatrix}K^{\mathrm{lat.},g/q}\\ J_{5}^{\mathrm{lat.},g/q}\end{pmatrix}$, (16)
+where we introduce the abbreviations $\mathcal{O}^{\mathrm{tree/lat.},{g/q}}\equiv\langle{g/q}|\mathcal{O}|{g/q}\rangle^{\mathrm{tree/lat.}}$ to simplify the notation.
+$|g\rangle$ or $|q\rangle$ denotes either gluon or quark external states with specific momentum and polarization.
+Note that $K^{\mathrm{tree},q}=J_{5}^{\mathrm{tree},g}=0$.
+Moreover, since only disconnected quark loops are involved, we have $J_{5}^{\mathrm{lat.},g}\approx 0$.
+Thus, the RI/MOM renormalization constants can be determined as:
+$Z_{11}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\frac{K^{\mathrm{tree,g}}}{K^{\mathrm{lat.,g}}},\quad\quad Z_{12}^{\mathrm{RI}}(\mu_{\rm RI}^{2})=-\frac{K^{\mathrm{lat.,q}}}{J_{5}^{\mathrm{lat.,q}}}Z_{11}^{\mathrm{RI}},$
+$Z_{21}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $0,\quad\quad\quad\quad\quad Z_{22}^{\mathrm{RI}}(\mu_{\rm RI}^{2})=\frac{J_{5}^{\mathrm{tree,q}}}{J_{5}^{\mathrm{lat.,q}}},$ (17)
+where $\mu_{\rm RI}$ denotes the renormalization scale in the RI/MOM scheme.
+With the above approximation, $Z_{22}^{\mathrm{RI}}$ reduces to the axial-vector current renormalization constant $Z_{A}^{\mathrm{RI}}$ in Ref.
+Hu and others (2024).
 In the spirit of $\mathrm{RI/MOM}$, the finite matching coefficients $R_{ij}$ can be defined in terms of the operator renormalization constants as follows:
-$\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}\equiv\begin{pmatrix}\bar{Z}^{\mathrm{DR}}_{11}&\bar{Z}^{\mathrm{DR}}_{12}\\ \bar{Z}^{\mathrm{DR}}_{21}&\bar{Z}^{\mathrm{DR}}_{22}\end{pmatrix}\cdot\begin{pmatrix}Z^{\mathrm{RI,\,DR}}_{11}&Z^{\mathrm{RI,\,DR}}_{12}\\ Z^{\mathrm{RI,\,DR}}_{21}&Z^{\mathrm{RI,\,DR}}_{22}\end{pmatrix}^{-1}\Big|_{\epsilon\rightarrow 0}=\begin{pmatrix}\bar{Z}^{\mathrm{lat}}_{11}&\bar{Z}^{\mathrm{lat}}_{12}\\ \bar{Z}^{\mathrm{lat}}_{21}&\bar{Z}^{\mathrm{lat}}_{22}\end{pmatrix}\cdot\begin{pmatrix}Z^{\mathrm{RI,\,lat}}_{11}&Z^{\mathrm{RI,\,lat}}_{12}\\ Z^{\mathrm{RI,\,lat}}_{21}&Z^{\mathrm{RI,\,lat}}_{22}\end{pmatrix}^{-1}\,.$ (27)
-Consequently, the $R_{ij}$ determined using the first equality $\bar{Z}^{\mathrm{DR}}\cdot\big(Z^{\mathrm{RI,\,DR}}\big)^{-1}\big|_{\epsilon\rightarrow 0}$ with DR can be employed, together with the input $Z^{\mathrm{RI,\,lat}}_{ij}\equiv Z^{\mathrm{RI}}_{ij}$ computed in (Total Gluon Helicity Contribution to Proton Spin from Lattice QCD), to obtain the values for the $\overline{\mathrm{MS}}$-renormalization constants $\bar{Z}^{\mathrm{lat}}_{ij}$ of the operator $K^{\mu}$ and $J_{5}^{\mu}$ but with UV-divergences regularized using Lattice, which can then be applied directly to the bare operator matrix elements computed in Lattice to obtain the $\overline{\mathrm{MS}}$-renormalized results.
-By virtue of (3), we can further derive the following shortcut expressions for $R_{ij}$:
+$\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}\equiv\begin{pmatrix}\bar{Z}^{\mathrm{DR}}_{11}&\bar{Z}^{\mathrm{DR}}_{12}\\ \bar{Z}^{\mathrm{DR}}_{21}&\bar{Z}^{\mathrm{DR}}_{22}\end{pmatrix}\cdot\begin{pmatrix}Z^{\mathrm{RI,\,DR}}_{11}&Z^{\mathrm{RI,\,DR}}_{12}\\ Z^{\mathrm{RI,\,DR}}_{21}&Z^{\mathrm{RI,\,DR}}_{22}\end{pmatrix}^{-1}\Big|_{\epsilon\rightarrow 0}=\begin{pmatrix}\bar{Z}^{\mathrm{lat}.}_{11}&\bar{Z}^{\mathrm{lat}.}_{12}\\ \bar{Z}^{\mathrm{lat}.}_{21}&\bar{Z}^{\mathrm{lat}.}_{22}\end{pmatrix}\cdot\begin{pmatrix}Z^{\mathrm{RI,\,lat}}_{11}&Z^{\mathrm{RI,\,lat}}_{12}\\ Z^{\mathrm{RI,\,lat}}_{21}&Z^{\mathrm{RI,\,lat}}_{22}\end{pmatrix}^{-1}\,.$ (18)
+Consequently, the $R_{ij}$ determined using the first equality $\bar{Z}^{\mathrm{DR}}\cdot\big(Z^{\mathrm{RI,\,DR}}\big)^{-1}\big|_{\epsilon\rightarrow 0}$ with DR can be employed, together with the input $Z^{\mathrm{RI,\,lat}}_{ij}\equiv Z^{\mathrm{RI}}_{ij}$ computed in Eq.
+(A.1), to obtain the values for the $\overline{\mathrm{MS}}$-renormalization constants $\bar{Z}^{\mathrm{lat}.}_{ij}$ of the operator $K^{\mu}$ and $J_{5}^{\mu}$ but with UV-divergences regularized by the lattice, which can then be applied directly to the bare operator matrix elements computed on the lattice to obtain the $\overline{\mathrm{MS}}$-renormalized results.
+By virtue of Eq.
+(16), we can further derive the following shortcut expressions for $R_{ij}$:
 $\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}=\begin{pmatrix}\frac{\overline{K}_{gg}}{K^{\mathrm{RI}}_{gg}}&\frac{\overline{K}_{\bar{q}q}}{J^{\mathrm{RI}}_{\bar{q}q}}\\ \frac{\overline{J}_{gg}}{K^{\mathrm{RI}}_{gg}}&\frac{\overline{J}_{\bar{q}q}}{J^{\mathrm{RI}}_{\bar{q}q}}\end{pmatrix}$ (19)
 where $\overline{K}(\overline{J})_{gg(\bar{q}q)}$ denote the Lorentz-invariant light-cone form-factors or matrix elements of the $\overline{\mathrm{MS}}$-renormalized $K(J)$-current operator but with $\mathrm{RI/MOM}$-renormalized external $g(q)$-fields.
-Note that $K^{\mathrm{RI}}_{gg}$ and $J^{\mathrm{RI}}_{\bar{q}q}$ equal to, respectively, the tree-level values of the aforementioned matrix elements of $K(J)$-current operator as a result of the $\mathrm{RI/MOM}$ renormalization condition (3).
+Note that $K^{\mathrm{RI}}_{gg}$ and $J^{\mathrm{RI}}_{\bar{q}q}$ are equal to, respectively, the tree-level values of the aforementioned matrix elements of $K(J)$-current operator as a result of the $\mathrm{RI/MOM}$ renormalization condition Eq.
+(16).
 Extra care is required in the calculation of the dimensionally-regularized bare matrix elements and the subsequent extraction of renormalization constants entering the r.h.s.
-of (27) and (28), because of the notorious issue in treating $\gamma_{5}$, and to less extent $\epsilon^{\mu\nu\rho\sigma}$ in DR, especially so in the present case due to the involvement of axial anomaly.
+of Eq.(18) and Eq.(19), because of the notorious issue in treating $\gamma_{5}$, and to a lesser extent $\epsilon^{\mu\nu\rho\sigma}$ in DR, especially in the present case due to the involvement of the axial anomaly.
 In view of the applications related to the study of polarized structure functions of the proton, we choose a (modified) $\overline{\mathrm{MS}}$ prescription where the gauge invariance and Adler-Bell-Jackiw equation are preserved (hence without chiral symmetry).
-So far, there is no practical prescription known to the authors to apply consistently an anticommuting $\gamma_{5}$ in DR for computing loop amplitudes with (closed) fermion chains with an odd number of $\gamma_{5}$, especially when the amplitudes are anomalous Chen:2023lus; Chen:2024hlv.
-On the other hand, the so-called Larin’s prescription Larin:1993tq of $\gamma_{5}$, where $\gamma_{5}$ in DR is treated non-anticommuting tHooft:1972tcz; Breitenlohner:1977hr; Breitenlohner:1975hg; Breitenlohner:1976te while the spacetime-metric tensors resulting from contracting a pair of $\epsilon^{\mu\nu\rho\sigma}$ are all in D-dimensions, is particularly suited for dealing with the QCD corrections at hand: both gauge invariance and the Adler-Bell-Jackiw equation are maintained and it is technically convenient to apply at high loop orders in QCD.
-With this set-up, we have derived the perturbative expressions for the bare matrix elements involved in (28) up to three loops, truncated to high orders in $\epsilon$ sufficient for obtaining their renormalized counterparts to $\mathcal{O}(\epsilon^{0})$; they are subsequently employed to extract the perturbative results for the finite $R_{ij}$ according to both (28) and (27), with perfect agreement found between the two.
+So far, there is no practical prescription known to the authors to apply consistently an anticommuting $\gamma_{5}$ in DR for computing loop amplitudes with (closed) fermion chains with an odd number of $\gamma_{5}$, especially when the amplitudes are anomalous Chen (2023, 2025).
+On the other hand, the so-called Larin’s prescription Larin (1993) of $\gamma_{5}$, where $\gamma_{5}$ in DR is treated non-anticommuting ’t Hooft and Veltman (1972); Breitenlohner and Maison (1977a, b, c) while the spacetime-metric tensors resulting from contracting a pair of $\epsilon^{\mu\nu\rho\sigma}$ are all in D-dimensions, is particularly suited for dealing with the QCD corrections at hand: both gauge invariance and the Adler-Bell-Jackiw equation are maintained and it is technically convenient to apply at high loop orders in QCD.
+With this set-up, we have derived the perturbative expressions for the bare matrix elements involved in Eq.(19) up to three loops, truncated to high orders in $\epsilon$ sufficient for obtaining their renormalized counterparts to $\mathcal{O}(\epsilon^{0})$; they are subsequently employed to extract the perturbative results for finite $R_{ij}$ according to both Eq.(19) and Eq.(18), with perfect agreement found between the two.
 (For more technical details of the computational workflow and evaluation of loop integrals, we kindly refer to Refs.
-Ahmed:2021spj; Chen:2021gxv.) For the sake of references, we document below the perturbative results for the finite matching coefficients $R_{ij}$ up to three loops in Landau gauge.
+Ahmed et al.
+(2021); Chen and Czakon (2022a).) For the sake of references, we document below the perturbative results for the finite matching coefficients $R_{ij}$ up to three loops in Landau gauge.
 We begin with $R_{11}$ that was utilized in the present work, which reads
@@
 The dependence of $R_{ij}$ on ${\mu}^{2}$ exhibited in the above perturbative expressions are governed by their RG equations.
-Following from the definition (27) the RG equations for the matching coefficients $R_{ij}$ read
+Following from the definition Eq.
+(18) the RG equations for the matching coefficients $R_{ij}$ read
 ${\mu}^{2}\,\frac{\mathrm{d}}{\mathrm{d}\,{\mu}^{2}}\,\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}=\begin{pmatrix}\gamma_{11}&\gamma_{12}\\ 0&\gamma_{22}\end{pmatrix}\cdot\begin{pmatrix}R_{11}&R_{12}\\ R_{21}&R_{22}\end{pmatrix}$, (24)
-where the matrix of anomalous dimensions in r.h.s.
-is fully determined by the (proper or modified) $\overline{\mathrm{MS}}$ renormalization constants in (27) in Larin’s prescription Larin:1993tq; these anomalous dimensions observe the relations Breitenlohner:1983pi; Ahmed:2021spj; Luscher:2021bog $\gamma_{11}=-\frac{\mathrm{d}\ln a_{s}}{\mathrm{d}\ln{\mu}^{2}}=-\beta$ and $\gamma_{22}=-2a_{s}\,n_{f}\,\gamma_{12}$ known now to four Chen:2021gxv and five loops Chen:2022lun
-With the explicit expression (29) for $R_{11}$, it is interesting to observe that in the limit of large $\beta_{0}$ or $n_{f}$, the leading terms in powers of $\mathcal{O}(a^{n}_{s}\,\beta^{n}_{0})$ exhibit the following neat pattern:
+where the matrix of anomalous dimensions on the r.h.s.
+is fully determined by the (proper or modified) $\overline{\mathrm{MS}}$ renormalization constants in Eq.
+(18) in Larin’s prescription Larin (1993); these anomalous dimensions observe the relations Breitenlohner et al.
+(1984); Ahmed et al.
+(2021); Lüscher and Weisz (2021) $\gamma_{11}=-\frac{\mathrm{d}\ln a_{s}}{\mathrm{d}\ln{\mu}^{2}}=-\beta$ and $\gamma_{22}=-2a_{s}\,n_{f}\,\gamma_{12}$ known now to four Chen and Czakon (2022a) and five loops Chen and Czakon (2022b)
+With the explicit expression Eq.
+(20) for $R_{11}$, it is interesting to observe that in the limit of large $\beta_{0}$ or $n_{f}$, the leading terms in powers of $\mathcal{O}(a^{n}_{s}\,\beta^{n}_{0})$ exhibit the following neat pattern:
 $R_{11}\Big|_{\text{large $\beta_{0}$}}\approx 1\,+\,a_{s}\,\beta_{0}\,\big(L_{\mu}+\frac{5}{3}\big)\,+\,a_{s}^{2}\,\beta_{0}^{2}\,\big(L_{\mu}+\frac{5}{3}\big)^{2}\,+\,a_{s}^{3}\,\beta_{0}^{3}\,\big(L_{\mu}+\frac{5}{3}\big)^{3}\,+\,\mathcal{O}(a_{s}^{4})\,$, (25)
@@
 Refer to caption Refer to caption
-Figure 7: Three-loop (left panel) and two-loop (right panel) matching coefficients in the $\mathrm{RI/MOM}$ scale at ${\mu}^{2}=10\ \mathrm{GeV}^{2}$.
+Figure 4: Matching coefficients in the $\mathrm{RI/MOM}$ scale at ${\mu}^{2}=10\ \mathrm{GeV}^{2}$.
+The matching coefficients $R_{ij}$ (two- and three-loop, in light and dark color, respectively) are shown in the left panel; the right panel specifically presents $R_{11}$ from one to three loops.
 In order to combine with lattice data, we change the horizontal axis to the dimensionless $\mathrm{RI/MOM}$ scale $a^{2}p^{2}$.
-Figure 8: One to three-loop matching coefficients $R_{11}$ in the $\mathrm{RI/MOM}$ scale at ${\mu}^{2}=10\ \mathrm{GeV}^{2}$.
-In order to combine with lattice data, we change the horizontal axis to the dimensionless $\mathrm{RI/MOM}$ scale $a^{2}p^{2}$.
-Table 2: Raw data employed for the global fitting corresponding to (16).
-The total uncertainty in the renormalization constants (for both the two- and three-loop matching) comprises two components: the statistical error and systematic error, which the systematic error given by difference between the center value of RG-evolved result and fixed-order result.
-$a$ (fm) Bare matrix elements $\bar{Z}_{11}^{\mathrm{lat}}$
-$p(\mathrm{GeV})$ $\langle K^{t}\rangle/(2E)$ $\langle K^{i}\rangle/(2E)$ $2\mathrm{-loop}$ $3\mathrm{-loop}$
+Furthermore, motivated by the features observed in Fig.
+4, it can be demonstrated that the off-diagonal elements of the matching coefficient are negligible.
+This leads to a significant simplification of the matching relations.
+Starting from simplified RI/MOM-renormalized gluon helicities $\Delta G$ and its $\overline{\mathrm{MS}}$ matching, namely
+$\Delta G^{\mathrm{RI}}$ $=$ $Z_{11}^{\mathrm{RI}}\big(\Delta G^{\mathrm{B.}}-\frac{K^{\mathrm{lat.,q}}}{J_{5}^{\mathrm{lat.,q}}}\Delta\Sigma^{\mathrm{B.}}\big),$
+$\Delta G^{\overline{\mathrm{MS}}}$ $=$ $R_{11}\Delta G^{\mathrm{RI}}+R_{12}\Delta\Sigma^{\mathrm{RI}},$ (26)
+we can obtain the exact formulations for gluon helicity in the $\overline{\mathrm{MS}}$ scheme
+$\Delta G^{\overline{\mathrm{MS}}}(a^{2}p^{2},{\mu}^{2})$ $=$ $R_{11}(a^{2}p^{2},{\mu}^{2})Z_{11}^{\mathrm{RI}}(a^{2}p^{2})\big(\Delta G^{\mathrm{B.}}-\frac{K^{\mathrm{lat.,q}}}{J_{5}^{\mathrm{lat.,q}}}(a^{2}p^{2})\Delta\Sigma^{\mathrm{B.}}\big)$ (27)
+$+R_{12}(a^{2}p^{2},{\mu}^{2})Z_{A}^{\mathrm{RI}}(a^{2}p^{2})\Delta\Sigma^{\mathrm{B.}},$
+where $Z_{11}^{\mathrm{RI}}=K^{\mathrm{tree,g}}/K^{\mathrm{lat.,g}}$ is the RI/MOM diagonal renormalization constant of gluon helicity, and we introduce the abbreviations $\mathcal{O}^{\mathrm{tree/lat.},{g/q}}\equiv\langle{g/q}|\mathcal{O}|{g/q}\rangle^{\mathrm{tree/lat.}}$ to simplify the notation.
+In addition, we express the dependence on the squared renormalization scale $\mu_{\rm RI}^{2}$ of the $\mathrm{RI/MOM}$ scheme in terms of the dimensionless quantity $a^{2}p^{2}$, and the same below.
+We now implement a three-loop approximation for the ratio term $K^{\mathrm{lat,q}}/{J_{5}^{\mathrm{lat,q}}}$.
+Noting that $R_{12}$ can be determined from the topological current $K^{\mu}$ for quark external states while $R_{22}\approx 1$ is fixed by the axial vector for quark external states.
+Finally we approximate
+$\big(\bar{Z}_{11}^{\mathrm{lat}.}R_{12}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2})$ $\approx$ $\big(\bar{Z}_{11}^{\mathrm{lat}.}\frac{R_{12}}{R_{22}}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2})$ (28)
+$=$ $\big(\bar{Z}_{11}^{\mathrm{lat}.}\frac{K^{3\mathrm{-loop,q}}}{J_{5}^{3\mathrm{-loop,q}}}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2})$
+$\approx$ $\big(\bar{Z}_{11}^{\mathrm{lat}.}\frac{K^{\mathrm{lat,q}}}{J_{5}^{\mathrm{lat,q}}}\big)(a^{2}p^{2}\rightarrow 0,{\mu}^{2}).$
+where we used a redefined expression $\bar{Z}_{11}^{\mathrm{lat}.}\equiv R_{11}Z_{11}^{\mathrm{RI}}$.
+Furthermore, owing to the protection by chiral symmetry, the matching coefficient for the axial-vector current remains unity up to two loops, i.e., $\bar{Z}_{A}^{\mathrm{lat}}\equiv Z_{A}^{\overline{\mathrm{MS}}}\approx Z_{A}^{\mathrm{RI}}$.
+This convention will be consistently maintained throughout the subsequent discussion.
+This approximation enables the following simplified expression for gluon helicity at three-loop order:
+$\Delta G^{\overline{\mathrm{MS}}}$ $\overset{3\mathrm{-loop}}{\approx}$ $\bar{Z}_{11}^{\mathrm{lat}.}\big(\Delta G^{\mathrm{B.}}-R_{12}\Delta\Sigma^{\mathrm{B}.}\big)+R_{12}\bar{Z}_{A}^{\mathrm{lat}.}\Delta\Sigma^{\mathrm{B.}}$ (29)
+$=$ $\bar{Z}_{11}^{\mathrm{lat}.}\Delta G^{\mathrm{B.}}+(\bar{Z}_{A}^{\mathrm{lat}.}-\bar{Z}_{11}^{\mathrm{lat.}})R_{12}\Delta\Sigma^{\mathrm{B.}}$
+$\approx$ $\bar{Z}_{11}^{\mathrm{lat}.}\Delta G^{\mathrm{B.}},$
+where the final reduction is justified based on the hierarchy $\bar{Z}_{11}^{\mathrm{lat}.}\gg(\bar{Z}_{A}^{\mathrm{lat}.}-\bar{Z}_{11}^{\mathrm{lat}.})R_{12}$, and by the fact that $\Delta G^{\mathrm{B.}}$ is comparable in magnitude to $\Delta\Sigma^{\mathrm{B.}}$.
+The above arguments are all self-evident in the lattie QCD calculations.
+The derivation for the expression $\Delta\Sigma^{\overline{\mathrm{MS}}}\approx\bar{Z}_{A}^{\mathrm{lat}.}\Delta\Sigma^{\mathrm{B.}}$ parallels the arguments used in the approximation above.
+A.2 B. Bare nucleon matrix elements with topological current in the moving frame
+The results in this work, are based on the $n_{f}=2+1$ flavor ensembles from the CLQCD collaboration using the tadpole improved tree level Symanzik (TITLS) gauge action and the tadpole improved tree level Clover (TITLC) fermion action Hu and others (2024).
+The TITLS gauge action is defined in the following,
+$S_{g}=\frac{1}{N_{c}}\mathrm{Re}\sum_{x,\mu<\nu}\mathrm{Tr}\left[1-\hat{\beta}\left(\mathcal{P}^{U}_{\mu,\nu}(x)+\frac{c_{1}\mathcal{R}^{U}_{\mu,\nu}(x)}{1-8c_{1}^{0}}\right)\right],$ (30)
+and
+$\mathcal{P}^{U}_{\mu,\nu}(x)$ $=U_{\mu}(x)U_{\nu}(x+a\hat{\mu})U^{\dagger}_{\mu}(x+a\hat{\nu})U^{\dagger}_{\nu}(x),$
+$\mathcal{R}^{U}_{\mu,\nu}(x)$ $=U_{\mu}(x)U_{\mu}(x+a\hat{\mu})U_{\nu}(x+2a\hat{\mu})U^{\dagger}_{\mu}(x+a\hat{\mu}+a\hat{\nu})U^{\dagger}_{\mu}(x+a\hat{\nu})U^{\dagger}_{\nu}(x),$
+$U_{\mu}(x)$ $=P\left[\mathrm{exp}\left({\rm i}g_{0}\int_{x}^{x+\hat{\mu}a}\mathrm{d}yA_{\mu}(y)\right)\right],$
+$\hat{\beta}=(1-8c_{1}^{0})\frac{6}{g_{0}^{2}u_{0}^{4}}\equiv 10/(g_{0}^{2}u_{0}^{4})$ with $c_{1}^{0}=-\frac{1}{12}$, $c_{1}=\frac{c_{1}^{0}}{u_{0}^{2}}$, $u_{0}=\langle\frac{\mathrm{Re}\mathrm{Tr}\sum_{x,\mu<\nu}\mathcal{P}^{U}_{\mu\nu}(x)}{6N_{c}\tilde{V}}\rangle^{1/4}$ is the tadpole improvement factor.
+The TITLC fermion action uses 1-step stout smeared link $V$ with smearing parameter $\rho=0.125$,
+$S_{q}(m)=\sum_{x,\mu=1,...,4,\eta=\pm}\bar{\psi}(x)\sum\frac{1+\eta\gamma_{\mu}}{2}V_{\eta\mu}(x)\psi(x{+\eta\hat{\mu}a})$
+$\quad+\sum_{x}\psi(x)\left[-(4+ma)\delta_{y,x}+c_{\rm sw}\sigma^{\mu\nu}g_{0}F_{\mu\nu}\right]\psi(x),$ (31)
+where $c_{\rm sw}=\frac{1}{v^{3}_{0}}$ with $v_{0}=\langle\frac{\mathrm{Re}\mathrm{Tr}\sum_{x,\mu<\nu}\mathcal{P}^{V}_{\mu\nu}(x)}{6N_{c}\tilde{V}}\rangle^{1/4}$, and the field strength tensor $F_{\mu\nu}(x)$ is defined through the following discretized formulations
+$F_{\mu\nu}(x)$ $=$ $\frac{i}{8g_{0}a^{2}}(\mathcal{P}_{\mu,\nu}^{U}(x)-\mathcal{P}_{\nu,\mu}^{U}(x)+\mathcal{P}_{\nu,-\mu}^{U}(x)-\mathcal{P}_{-\mu,\nu}^{U}(x)$ (32)
+$+\mathcal{P}_{-\mu,-\nu}^{U}(x)-\mathcal{P}_{-\nu,-\mu}^{U}(x)+\mathcal{P}_{-\nu,\mu}^{U}(x)-\mathcal{P}_{\mu,-\nu}^{U}(x)).$
+The detailed information for the simulation is given in Table 2.
+For all ensembles except C48P23, the quark propagators are computed for all time slices using the distillation smearing method Peardon et al.
+(2009), 1-step stout smearing is applied to the links in the Laplacian operator (with smearing parameter of $\rho=0.125$), and number of eigenvectors $N_{\mathrm{ev}}$ in the smearing operator is 100.
+For the proton external states with momentum greater than $0.5$ GeV, momentum smearing Bali et al.
+(2016); Egerer et al.
+(2021) is further implemented using smearing parameter $\textbf{k}=(0,0,2)$.
+The C48P23 ensemble is intended specifically for renormalization calculations; therefore, neither distillation nor momentum smearing is employed.
+Symbol $\hat{\beta}$ $a$ (fm) $\tilde{L}^{3}\times\tilde{T}$ $m_{\pi}$ (MeV) $N_{\mathrm{cfg}}$ $\rho$ $N_{ev}$ k
+C24P29 6.200 0.10524(05)(62) $24^{3}\times 72$ 292.3(1.0) 780 0.125 100 $(0,0,2)$
+C48P23 $48^{3}\times 96$ 224.1(1.2) 400 0.125 - -
+E32P29 6.308 0.08973(20)(53) $32^{3}\times 64$ 287.3(2.5) 890 0.125 100 $(0,0,2)$
+F32P30 6.410 0.07753(03)(45) $32^{3}\times 96$ 300.4(1.2) 800 0.125 100 $(0,0,2)$
+Table 2: Lattice setup used for the simulation, which including gauge coupling $\hat{\beta}=10/(g_{0}^{2}u_{0}^{4})$, lattice spacing $a$, dimensionless lattice size $\tilde{L}^{3}\times\tilde{T}$, corresponding pion mass $m_{\pi}$, number of configurations $N_{\mathrm{cfg}}$, 1-step stout smearing parameter $\rho$, number of distillation eigenvectors $N_{\mathrm{ev}}$ and momentum smearing parameter k.
+In lattice computations of the topological current
+$K^{\mu}(x)=\epsilon^{\mu\nu\rho\sigma}\text{Tr}[A_{\nu}F_{\rho\sigma}-\frac{2}{3}ig_{0}A_{\nu}A_{\rho}A_{\sigma}](x),$ (33)
+the fundamental components requiring evaluation are the gauge potential $A_{\mu}(x)$ and the field strength tensor $F_{\mu\nu}(x)$, where $A_{\mu}(x)$ is defined through the following discretized formulations
+$A_{\mu}(x+\hat{\mu}/2)=\left[\frac{U_{\mu}(x)-U_{\mu}^{\dagger}(x)}{2ig_{0}a}\right]_{\mathrm{traceless}.}.$ (34)
+While an alternative full-lattice definition $\tilde{A}_{\mu}(x)=[A_{\mu}(x+\hat{\mu}/2)+A_{\mu}(x-\hat{\mu}/2)]/2$ exists, the half-lattice and full-lattice forms of the gauge potential differ only in their discretization errors.
+We therefore use the half-lattice definition in all following calculations.
+Based on definition of the local topological current $K^{\mu}$ in Eq.
+(33), its temporal and spatial components read explicitly:
+$K_{i}$ $\equiv$ $K^{E\times A}_{i}+K_{i}^{A_{t}B}+K_{i}^{3A}=2\text{Tr}[(E\times A)_{i}]-2\text{Tr}[A_{t}B_{i}]+\frac{2}{3}ig_{0}\epsilon^{ijk}\mathrm{Tr}[A_{t}[A_{j},A_{k}]]$
+$K_{t}$ $\equiv$ $K_{t}^{B\cdot A}+K_{t}^{3A}=2\text{Tr}[B\cdot A]-\frac{2}{3}ig_{0}\epsilon^{ijk}\mathrm{Tr}[A_{i}A_{j}A_{k}],$ (35)
+where $E_{i}=F_{ti}$ and $B_{i}=F_{jk}$.
+Note that the component expression of $K^{\mu}$ is defined in Euclidean space.
+Due to signal-to-noise ratio considerations, 5-HYP smearing is applied to the topological current throughout this simulation.
+Following Ref. Pang et al. (2024), the total gluon helicity $\Delta G$ is obtained from the matrix element of the local topological current,
+$\frac{\langle\mathrm{PS}_{\mathrm{Proj.i}}|\int d^{3}xK^{\mu}_{\mathrm{C.G.}}(x)|\mathrm{PS}_{\mathrm{Proj.i}}\rangle}{2S^{\mu}\langle\mathrm{PS}_{\mathrm{Unproj.}}|\mathrm{PS}_{\mathrm{Unproj.}}\rangle}|_{\mathrm{IMF}}=\Delta G.$ (36)
+[…]
+By defining the enhanced proton operator Zhang et al. (2025)
+$N(\vec{p},t)=\sum_{\vec{x}}e^{-i\vec{p}\cdot\vec{x}}P_{+}(u^{T}C\gamma_{5}\gamma_{t}du)(\vec{x},t),$ (37)
+the two-point correlation function (2pt) can be defined as
+$C_{2}(p,t_{f})=\sum_{t_{s}}\langle 0|N(p,t_{s}+t_{f})N^{\dagger}(p,t_{s})|0\rangle,$ (38)
+with source time slice $t_{s}$, difference between source and sink time slice $t_{f}$.
+Extracting the effective mass from two-point correlation functions is an essential procedure, as it isolates the ground-state contribution—a prerequisite for reliably determining baryon matrix elements.
+FIG.
+5 presents behavior of the proton effective mass for ensemble C24P29.
+Our analysis shows that while the distillation smearing technique alone provides well-optimized plateau regions for the effective mass at low momenta, this stability deteriorates significantly at higher momenta (e.g., p=1.96GeVp=1.96\ $\mathrm${GeV}).
+In comparison, implementation of distillation + momentum smearing scheme yields moderately improved plateau stability at large momenta relative to the distillation-only approach.
+Figure 5: Proton effective mass with different momentum for ensemble C24P29.
+The data points with error represent the distillation smearing scheme alone, while the error bands represent the distillation + momentum smearing.
+The red, green and blue data points (or bands) represent p=0,0.98,1.96GeVp=0,0.98,1.96\ $\mathrm${GeV}, respectively.
+In addition, the three-point correlation function (3pt) can be defined as
+$C_{3}(p,t_{f},t_{i})=\sum_{t_{s}}\langle 0|\Gamma_{i}N(p,t_{s}+t_{f})\big(\sum_{\vec{x}}K^{\mu}_{\mathrm{C.G.}}(x,t_{i})\big)N^{\dagger}(p,t_{s})|0\rangle,$ (39)
+with current insertion time slice $t_{i}$ and $\Gamma_{i}=\gamma_{i}\gamma_{5}$.
+Among Eq.
+(38) and Eq.
+(39), the topological current bare matrix elements (BMEs) for the momentum $p$ can be obtained through the following ratio
+$R_{K^{\mu}}(p,t_{f},t_{i})=\frac{C_{3}(p,t_{f},t_{i})}{C_{2}(p,t_{f})},$ (40)
+which is then fitted to the form
+$R_{K^{\mu}}(p,t_{f},t_{i})=\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}(p)+c_{1}e^{-\Delta E(t_{f}-t_{i})}+c_{2}e^{-\Delta Et_{i}},$ (41)
+where $\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}$ is defined as nuclear BME divided by the field operator’s normalization factor $2E$, and $c_{1},c_{2},\Delta E$ are fitting parameters.
+The matrix elements of the topological current can be effectively discussed by separating its temporal and spatial components.
+The temporal component $K_{t}$, dominated by the magnetic field strength term $K_{t}^{B\cdot A}$, yields matrix elements consistent with zero within uncertainties at zero momentum.
+These matrix elements then exhibit a gradual increase before saturating to a constant value at higher momenta.
+Conversely, the spatial components $K_{i}$, governed primarily by the electric field strength term $K_{i}^{E\times A}$, maintain approximately constant matrix elements within errors across all momenta studied.
+This behavior finds a natural interpretation within a point-like proton model, where $\langle K^{\mu}\rangle_{N}\propto S_{\mu}/E$: the charge density (related to $K_{t}$) is an intrinsic property, while the current density (related to $K_{i}$) increases from zero at rest toward a finite asymptotic value.
+These distinct momentum dependencies constitute a direct manifestation of Lorentz covariance in a given frame.
+Using the same smearing technique employed for the proton effective mass, we extract the BMEs of $K^{t/i}$.
+The fitted BMEs presented in FIG.
+6 validate this theoretical picture, thereby confirming the description outlined above and providing direct evidence for the relations discussed around Eq.
+(A.2).
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 6: Ratio $R_{K^{\mu}}(t_{f},t_{i})$ at $t_{f}\in[0.63,1.16]\ \mathrm{fm}$ and the BMEs $\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}$ (the gray error band) for proton external states with varying momenta for ensemble C24P29, utilizing the fitting procedure described in Eq.
+(41).
+The results are organized such that the first column corresponds to spatial components $K_{i}$, while the second column represents temporal component $K_{t}$.
+Rows one through three display results for proton momenta of p=0,0.98,1.96GeVp=0,0.98,1.96\ $\mathrm${GeV}, respectively.
+A.3 C. Non-perturbative renormalization with cluster decomposition error reduction (CEDR) and $a^{2}p^{2}$ correction
+Within the non-perturbative RI/MOM scheme, the topological current’s diagonal renormalization constant $Z_{11}^{\mathrm{RI}}=K^{\mathrm{tree,g}}/K^{\mathrm{lat.,g}}$ can be determined as follows
+$Z_{11}^{\mathrm{RI}}(\mu_{\rm RI}^{2})=\frac{Z_{g,\mathrm{diag/off}}(\mu_{\rm RI}^{2})}{\mathrm{Tr}[S_{g}^{-1}(p_{1})G_{K}(p_{1},p_{2})S_{g}^{-1}(p_{2})(K^{\mathrm{tree,g}})^{-1}]},$ (42)
+where the external momenta are chosen to satisfy $p_{1}=p_{2}=p$ and $\mu_{\rm RI}^{2}=p^{2}$ denotes the RI/MOM renormalization scale. The bare gluon propagator is given by
+$S_{g}^{\mu\nu}(p)=\mathrm{C_{V}}\langle\mathrm{Tr}[A^{\mu}(p)A^{\nu}(-p)]\rangle,$ (43)
+where $\mathrm{C_{V}}=2/[(N^{2}_{c}-1)\tilde{V}$], with $N_{c}=3$ representing the number of colors and $\tilde{V}=\tilde{L}^{3}\times\tilde{T}$ denoting the dimensionless lattice volume.
+Due to rotational symmetry breaking on the lattice, we can also define the gluon field renormalization constants through two distinct prescriptions
+$Z_{g,\mathrm{diag}}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\frac{\sum_{p_{\rho}=0}1}{\hat{p}^{2}\sum_{p_{\rho}=0}S_{g}^{\rho\rho}(p^{2})},$
+$Z_{g,\mathrm{off}}^{\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\frac{\sum_{p_{\rho}\neq 0,p_{\tau}\neq 0,\rho\neq\tau}1}{\hat{p}^{4}\sum_{p_{\rho}\neq 0,p_{\tau}\neq 0,\rho\neq\tau}\frac{S_{g}^{\rho\tau}(p^{2})}{\hat{p}_{\rho}\hat{p}_{\tau}}},$ (44)
+where $\hat{p}_{\mu}=\frac{2}{a}\sin{\frac{p_{\mu}a}{2}}$ represents the lattice momentum.
+In addition, the computation of $Z_{11}^{\mathrm{RI}}$ necessitates the evaluation of the bare Green’s function
+$G_{K}^{\mu\rho\nu}(p_{1},p_{2})=\sum_{x,y}e^{-i(p_{1}\cdot x-p_{2}\cdot y)}\langle A^{\mu}(x)K^{\rho}(0)A^{\nu}(y)\rangle,$ (45)
+both $A^{\mu}(x)$ and $K^{\mu}(x)$ are accessible to direct lattice computation via Eq. (34) and Eq. (33). Ultimately, from all the above formulas, the inverse of $Z_{11}^{\rm RI}$ then reads
+$Z_{11}^{-1,\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\sum_{\mu,\rho,\nu,\sigma}\frac{2\mathrm{C_{V}}\hat{p}^{2}G_{K}^{\mu\rho\nu}(p,p)}{i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}(S_{g}^{\mu\mu}+S_{g}^{\nu\nu})(p^{2})}|_{\mu_{\rm RI}^{2}=p^{2}}\text{ or}$
+$Z_{11}^{-1,\mathrm{RI}}(\mu_{\rm RI}^{2})$ $=$ $\sum_{\mu,\rho,\nu,\sigma}\frac{\mathrm{C_{V}}\hat{p}^{4}G_{K}^{\mu\rho\nu}(p,p)}{i\epsilon^{\rho\mu\sigma\nu}\hat{p}_{\sigma}\hat{p}_{\mu}\hat{p}_{\nu}S_{g}^{\mu\nu}(p^{2})}|_{\mu_{\rm RI}^{2}=p^{2}}.$ (46)
+After matching, the logarithmic divergences are canceled, leaving only a polynomial dependence on the dimensionless RI/MOM scale $a^{2}p^{2}$.
+Therefore, in the subsequent analysis, we employ the relation
+$\bar{Z}_{11}^{\mathrm{lat.}}(a^{2}p^{2},\mu^{2})=R_{11}(a^{2}p^{2},\mu^{2})Z_{11}^{\mathrm{RI}}(a^{2}p^{2}),$ (47)
+and set $\mu^{2}=10\ \mathrm{GeV}^{2}$, where $R_{11}$ and $Z_{11}^{\mathrm{RI}}$ are given by Eq. (20) and Eq. (A.3), respectively.
+Despite achieving a precision better than 1% in the determination of the gluon propagator (Eq.
+(43)), the calculation of the bare gluon Green’s function (Eq.
+(45)) remains subject to formidable numerical challenges.
+More concretely, the direct computation of Eq.
+(45) is plagued by substantial statistical uncertainties, making a stable determination of $Z_{11}^{\mathrm{RI}}$ (equivalently $\bar{Z}_{11}^{\mathrm{lat.}}$) at any scale viable only with a statistical sample exceeding hundreds of thousands of configurations.
+Applying the cluster decomposition error reduction (CDER) technique enables significant error suppression in $Z_{11}^{\mathrm{RI}}$ (or $\bar{Z}_{11}^{\mathrm{lat.}}$) calculations Liu et al.
+(2018); Yang et al.
+(2018a).
+The fundamental principle of cluster decomposition asserts that correlators exhibit exponential decay with increasing separation between operator insertions, implying that integration beyond the correlation length primarily contributes to statistical noise rather than meaningful signal.
+The CDER methodology addresses this by truncating the volume integral beyond a characteristic distance, thereby achieving an enhancement of the signal-to-noise ratio by a factor of $\sqrt{V}$, as demonstrated in Ref.
+Liu et al.
+(2018).
+The CDER calculation for Eq. (45) can be achieved through the following expression
+$G_{K}^{\mu\rho\nu,\mathrm{CDER}}(p,p)=\langle\int_{|r|<r_{1}}d^{4}r\int_{|r^{\prime}|<r_{2}}d^{4}r^{\prime}\int d^{4}xe^{ip\cdot r^{\prime}}K^{\rho}(x+r^{\prime})\mathrm{Tr}[A_{\mu}(x)A_{\nu}(x+r)]\rangle,$ (48)
+where we introduce two distinct cutoff parameters: $r_{1}$ between the gluon operator and one gauge field, and $r_{2}$ between the gauge fields within the gluon propagator.
+This implementation yields the modified correlator with appropriate spatial cutoffs.
+The CDER technique achieves computational efficiency of $\mathcal{O}(V\log V)$ through the repeated application of fast Fourier transforms (FFTs).
+In concrete terms, Eq.
+(48) can be realized through the strategy outlined below:
+1) Construct the spatially-restricted topological current operator: We compute $K^{\rho}_{r^{\prime}}(x)=\int_{|r^{\prime}|<r_{2}}d^{4}r^{\prime}K^{\rho}(x+r^{\prime})$ by first performing Fourier transforms (FT) of both $K^{\rho}(x)$ and the Heaviside function $f(x)=\theta(r_{2}-|x|)$, multiplying the transformed functions in momentum space, and subsequently applying anti-FT.
+2) Form the intermediate operator product: We calculate $B^{\rho}_{\mu,r^{\prime}}(x)=A_{\mu}(x)K^{\rho}_{r^{\prime}}(x)$ through direct multiplication in coordinate space.
+3) Apply cluster decomposition to the correlation function: For the expression $\int d^{4}xd^{4}ye^{ip\cdot(x-y)}B^{\rho}_{\mu,r^{\prime}}(x)A_{\nu}(y)$, we perform FT of both $A_{\nu}$ and $B^{\rho}_{\mu,r^{\prime}}$, apply anti-FT to the product $A_{\nu}(p)B^{\rho}_{\mu,r^{\prime}}(-p)$, implement an additional Heaviside function $g(x)=\theta(r_{1}-|x|)$ in coordinate space, and FT the resulting product.
+The selection of cutoff parameters $r_{1,2}$ in the CDER technique requires careful consideration, as inappropriate choices may induce periodic oscillations in the renormalization constants Liu et al.
+(2018).
+FIG.
+7 provides visual confirmation of this analysis under ensemble C48P23.
+When the cutoff parameter governing the gluon field-operator correlation ($r_{1}$) is set too small (top left panel), distinct periodic oscillations become apparent in the results.
+These oscillations progressively stabilize as $r_{1}$ increases.
+A similar pattern is observed for variations in the current cutoff parameter ($r_{2}$), though changes in oscillatory behavior are less pronounced without corresponding adjustments to $r_{1}$.
+Based on the above discussion, we employ the optimal cutoffs identified in the bottom right panel of FIG.
+7—specifically, $r_{1}=1.26\ \mathrm{fm}$ and $r_{2}=0.74\ \mathrm{fm}$—to perform the subtraction of discretization errors.
+This choice of cutoff parameters aligns with the observation in Ref.
+Yang et al.
+(2018a) that $r_{1}\approx 1.4r_{2}$ yields the optimal signal-to-noise performance.
+Refer to caption Refer to caption
+Refer to caption Refer to caption
+Figure 7: Performance of the renormalization constant $\bar{Z}_{11}^{\mathrm{lat.}}(a^{2}p^{2},{\mu}^{2}=10\ \mathrm{GeV}^{2})$ for various cutoff parameters $r_{1,2}$ under ensemble C48P23 using momentum modes $(p,p,p,p)$.
+As established, $\bar{Z}_{11}^{\mathrm{lat}.}$ carries a discretization error that depends on $a^{2}p^{2}$.
+However, as evidenced by Fig.
+7, within the effective window of $a^{2}p^{2}$—determined by considering both perturbative matching and the signal-to-noise ratio of the lattice data—this discretization error remains too large to permit a reliable fit.
+Consequently, we employ the method from Ref.
+Yang et al.
+(2018b) to suppress the $a^{2}p^{2}$ corrections.
+Specifically, for the 5-HYP smeared topological current, most of the $a^{2}p^{2}$ dependence in $\bar{Z}{11}^{\mathrm{lat}.}$ can be eliminated by constructing the following ratio:
+$\tilde{\bar{Z}}_{11}^{\mathrm{lat.}}(a^{2}p^{2},\mu^{2})=\bar{Z}_{11}^{\mathrm{lat.}}(a^{2}p^{2},\mu^{2})f(a^{2}p^{2}\rightarrow 0)/f(a^{2}p^{2}),\quad f(a^{2}p^{2})=\frac{\langle\mathrm{Tr}[A_{\mu}^{5\mathrm{-HYP}}(p)A_{\mu}^{5\mathrm{-HYP}}(-p)]\rangle}{\langle\mathrm{Tr}[A_{\mu}(p)A_{\mu}(-p)]\rangle},$ (49)
+where $A_{\mu}^{5\mathrm{-HYP}}$ is the 5-HYP-smeared gauge potential defined from the 5-HYP-smeared gauge link, and $f(a^{2}p^{2})$ is the smeared ratio of two gluon propagators (Eq.(43)) with and without HYP smearing.
+Fig.
+8 demonstrates the significance of the $a^{2}p^{2}$ correction.
+Prior to correction, the renormalization constant exhibits a severe dependence on $a^{2}p^{2}$, rendering the subtraction of discretization errors unpredictable.
+After applying the correction, although the signal-to-noise ratio deteriorates—primarily because the factor $f(a^{2}p^{2}\rightarrow 0)/f(a^{2}p^{2})$ itself becomes large at high $a^{2}p^{2}$—the result forms a plateau near unity.
+This behavior aligns with the physical expectation for a renormalization constant.
+Figure 8: Behavior of the renormalization constant under three-loop matching for ensemble C48P23, comparing results with (green) and without (red) the $a^{2}p^{2}$ correction from Eq.
+(49).
+The corresponding error bands represent fits that subtract discretization errors using an $a^{2}p^{2}$ polynomial: a third-order polynomial for the green band and a first-order polynomial for the red band.
+The fitted interval is $a^{2}p^{2}\in[2.6,7.1]$.
+The above analysis based for other ensembles (such as E32P29 and F32P30 in Table 2) give similar results.
+Thus, we have successfully determined the purely gluonic renormalization constant by employing the CDER technique combined with $a^{2}p^{2}$ correction.
+The CDER technique effectively suppresses the severe statistical noise inherent in direct calculations by truncating the volume integral beyond characteristic distances $r_{1}$ and $r_{2}$.
+Concurrently, the $a^{2}p^{2}$ correction significantly reduces the discretization errors in the renormalization constant, yielding reliable physical results.
+Together, these techniques allow us to determine this quantity with acceptable error—a task previously feasible only by accumulating enormous statistics.
+Building upon all the preceding analysis, we perform a combined extrapolation to the continuum limit and the IMF, namely
+$\langle K^{\mu}\rangle_{N}=\bar{Z}_{11}^{\mathrm{lat.}}\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}=\frac{S^{\mu}}{E}\big(\Delta G+\frac{c_{\mathrm{h.t.}}}{E^{2}}\big)+c_{a}a^{2},$ (50)
+where $c_{\mathrm{h.t.}}$ accounts for higher-twist contributions, and $c_{a}$ captures discretization effects.
+Since the proton momenta are not significantly larger than its mass, we incorporate the higher-twist contribution using an empirical parametrization form $1/E^{2}$ in the equation above, where $E=\sqrt{M^{2}+p^{2}}$ with the proton mass $M\approx 1\ \mathrm{GeV}$.
+To maintain clarity and avoid symbol redundancy, we adopt the notation $\bar{Z}_{11}^{\mathrm{lat}.}$ in place of $\tilde{\bar{Z}}_{11}^{\mathrm{lat}.}$ in the following discussion.
+This extrapolation utilizes data from three distinct lattice spacings and five proton momentum values, with the input raw data provided in Table 3.
+In addition, the primary sources of systematic uncertainty in our analysis are identified as follows:
+1) Including or not the zero-momentum data points in the extrapolation;
+2) Discretization errors associated with the finite lattice spacing: the central value difference between the zero momentum $\langle K_{i}\rangle_{N}$ on finest lattice spacing and that extrapolated to the continuum limit;
+3) The choice of the UV window in the RI/MOM scale or $a^{2}p^{2}$: the interval $\mu_{\mathrm{RI}}\in[3.0,5.0]\ \mathrm{GeV}$ has been extended to [3.0,6.0]GeV[3.0,6.0]\ $\mathrm${GeV} to estimate the corresponding uncertainty;
+4) Systematic errors associated with the perturbative inputs: estimated by varying from three-loop to two-loop accuracy.
+As illustrated in FIG.
+9, the joint extrapolation yields a gluon helicity contribution $\Delta G^{\overline{\mathrm{MS}},\ 10\ \mathrm{GeV}^{2}}=0.231(17)^{\mathrm{sta.}}(44)^{\mathrm{sym.}}$, which constitutes $46(9)\%$ of the proton’s total spin.
+$a$ (fm) $\langle K^{\mu}\rangle_{N}^{\mathrm{B.}}$ $\bar{Z}_{11}^{\mathrm{lat}.}$
+$p(\mathrm{GeV})$ $\langle K_{t}\rangle_{N}^{\mathrm{B.}}$ $\langle K_{i}\rangle_{N}^{\mathrm{B.}}$ $2\mathrm{-loop}$ $3\mathrm{-loop}$
 0.10524(05)(62) 0.00 0.006(05) 0.212(30) 0.91(09)(03) 1.09(12)(04)
@@
 1.50 0.173(30) 0.176(32)
+Table 3: Raw data employed for the global fitting corresponding to Eq.
+(50).
+The total uncertainty in the renormalization constants (for both the two- and three-loop matching) comprises two components: the statistical error and systematic error, which the systematic error given by difference between the center value of RG-evolved result and fixed-order result.
+Figure 9: The continuum and infinite momentum extrapolation of the total gluon helicity is performed using lattice spacings from 0.08fm0.08\ $\mathrm${fm} to 0.11fm0.11\ $\mathrm${fm} at the $\overline{\mathrm{MS}}$ scale $\mu^{2}=10\ \mathrm{GeV}^{2}$.
+The extrapolation employs the functional form given in Eq.
+(50), with the raw input data provided in Table 3.
+The uncertainties are shown as separate bands: statistical uncertainty (dark gray) and all uncertainties that arise during our analysis (light gray).